### v1

# Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling

Abstract
Accurate prediction of clinical outcomes using Electronic Health Records (EHRs) is critical for early intervention, efficient resource allocation, and improved patient care. EHRs contain multimodal data, including both structured data and unstructured clinical notes that provide rich, context-specific information. In this work, we introduce a unified framework that seamlessly integrates these diverse modalities, leveraging all relevant available information through a two-stage architecture for clinical risk prediction. In the first stage, a fine-tuned Large Language Model (LLM) extracts crucial, task-relevant information from clinical notes, which is enhanced by graph-based retrieval of external domain knowledge from sources such as a medical corpus like PubMed, grounding the LLM’s understanding. The second stage combines both unstructured representations and features derived from the structured data to generate the final predictions. This approach supports a wide range of clinical tasks. Here, we demonstrate its effectiveness on 30-day readmission and in-hospital mortality prediction. Experimental results show that our framework achieves strong performance, with AUC scores of and , respectively, despite these tasks involving severely imbalanced datasets, with positive rates ranging from approximately to . Moreover, it outperforms all existing baselines and clinical practices, including established risk scoring systems. To the best of our knowledge, this is one of the first frameworks for healthcare prediction which enhances the power of an LLM-based graph-guided knowledge retrieval method by combining it with structured data for improved clinical outcome prediction.
1 Introduction
Appropriate use of clinical prediction tools for early identification of high-risk patients for different conditions allows for clinical decision making, timely interventions, escalation of care, intensive monitoring, and identification of gaps in outpatient management. e.g., (Cai et al. 2016; Kong, Lin, and Hu 2020; Brajer et al. 2020; Kansagara et al. 2011; Mahmoudi et al. 2020). For instance, readmission within a short period is a priority under many regulatory frameworks and value-based care models, where high readmission rates may lead to financial penalties; e.g., (Upadhyay, Stephenson, and Smith 2019; Clement et al. 2017). Therefore, good models of short-term risk prediction (e.g., 30 days) can guide targeted interventions such as more detailed discharge instructions, closer post-discharge monitoring, or referrals to transitional care programs. While traditional prediction tools relied on logistic regression for risk assessment, machine learning methods are increasingly applied to clinical prediction tasks.
There has been a lot of work on developing diverse kinds of machine learning methods for such risk prediction questions using Electronic Health Record (EHR) data, which contain rich information on patient health, e.g., (Cai et al. 2016; Jiang et al. 2025; Qiu et al. 2025; Yang et al. 2023; Brisimi et al. 2018; Bellot and van der Schaar 2019; Cui et al. 2025b; Arsalan 2025; Yu et al. 2024; Wu, Xiao, and Sun 2023; Zhu et al. 2024). Most of this work focuses on structured EHR data, which spans various types and representations (Hartvigsen et al. 2018), including admission/discharge information, procedures and interventions, medications, lab orders and results, billing codes (e.g., ICD, CPT), physiological time-series (e.g., vital signs), imaging data, and clinical documentation such as progress notes or discharge summaries. When used for machine learning approaches to clinical prediction, unstructured data from clinical notes has been used in a fairly simple manner, such as bag of words, TF-IDF (Wu and Zhang 2021; Huang, Osorio, and Sy 2019) to facilitate the use of conventional machine learning methods for clinical tasks. Clinical notes are complex and poorly structured, which limits their utility in clinical informatics tasks, even when using advanced natural language processing (NLP) techniques. While large language models (LLMs) offer a powerful means to process such notes, especially when combined with large biomedical datasets to capture richer semantics beyond keywords and embeddings (Shi et al. 2024; Li et al. 2024; Singhal et al. 2023; Jiang et al. 2025), they still face significant limitations. Hallucinations, factual inaccuracies, and a lack of domain grounding make their use in clinical tasks challenging. Even standard methods such as retrieval-augmented generation (RAG) have produced inaccurate predictions in clinical settings (Shi et al. 2024; Jiang et al. 2025). For instance, models such as Med-PaLM (Singhal et al. 2023) and EHR-Chain-of-Thought (Jiang et al. 2025) exhibit strong language generation capabilities but frequently misinterpret similar-sounding medical terms. Similarly, RAG-based systems like EHRAgent (Shi et al. 2024) struggle when the retrieved context fails to align with a patient’s history. Recent approaches have attempted to enhance LLMs with structured knowledge via graph-based retrieval (e.g., GraphRAG), but their performance remains limited due to a lack of explicit reasoning. Efforts such as MedRetriever (Ye et al. 2021), GraphCare (Jiang et al. 2024), RAM-EHR (Xu et al. 2024), EHR-KnowGen (Niu et al. 2024), and EMERGE (Zhu et al. 2024) inject knowledge into patient representations, yet still fall short of leveraging the full reasoning potential of LLMs. In recent work, Jiang et al. (Jiang et al. 2025) developed KARE, a graphRAG and context augmentation approach, for clinical prediction tasks, which address many challenges associated with using LLMs for clinical tasks on the MIMIC dataset. However, the performance of all prior methods remains limited, and these methods are not usable in clinical decision making.
Challenges. Clinical tasks using EHRs present the following non-trivial challenges: (1) Multi-modal Information: Clinical records in EHRs include both structured and unstructured data. How can we effectively handle both modalities within a single processing pipeline? (2) Long Context: Clinical documents are typically verbose and repetitive. How can we extract relevant information efficiently while avoiding distractions from redundant content? (3) Specialized Medical Domains: EHRs involve domain-specific terminology, such as drugs and treatments. How can we ensure accurate interpretation of such specialized medical information? (4) Highly imbalanced data: Clinical datasets are typically very imbalanced, e.g., about 4% positive for 30-day readmission and about 13% positive for mortality, which has a severe impact on the performance of most ML methods. We will address these questions in the framework Section 4.
Our contributions.
1. We develop a novel framework, KAMELEON (Knowledge-Augmented Multimodal EHR LEarning for Outcome predictioN)
that integrates multimodal EHR data (including structured clinical components and unstructured physician notes), and external biomedical knowledge.
KAMELEON consists of two components (Figure 1):
(1) a graphRAG and knowledge-augmented reasoning approach, which builds on (Jiang et al. 2025), and outputs a prediction and reasoning; and
(2) a structured machine learning module which uses tabular and time-series datasets in the EHR, along with the outputs of the LLM based component.
2. We demonstrate the effectiveness of KAMELEON for two clinical tasks: 30-day readmission risk and mortality prediction, which have been studied extensively, both using MIMIC-III and other private EHRs from specific hospitals.
We compare the performance with a number of structured ML and LLM baselines, with respect to multiple metrics.
KAMELEON achieves significantly better performance for these tasks than all prior work on MIMIC-III datasets.
We observe a relative improvement in the AUROC of approximately for the 30-day clinical readmission prediction task, and around for the mortality prediction task.
Similarly, we achieve a relative AUROC improvement of for 30-day readmission and for in-hospital mortality over the best unstructured LLM baseline (Llama3-Med42-8B). Our work outperforms clinical score methods as well.
The only other prior work which has similar performance for 30-day readmission (Bopche et al. 2024) is on a Norwegian EHR dataset, which is significantly less imbalanced ( readmission positive rate, instead of in the case of MIMIC-III).
In summary, KAMELEON is the first systematic framework to enhance the power of LLMs for healthcare prediction tasks by using graph-guided knowledge retrieval with structured machine learning methods; we expect this framework to be easily applicable to other clinical questions beyond the ones we study here.
Note: Many details of the methodology and additional results are included in the Appendix for brevity.
2 Related Works
Clinical Prediction using Statistical and Machine Learning Methods: Structured data collected from electronic health records (EHRs) contain valuable information about patients and can be instrumental in capturing patient states and predicting medical outcomes (e.g., mortality) (Jiang et al. 2025; Qiu et al. 2025; Yang et al. 2023). To leverage these data, traditional machine learning models—such as linear regression (Brisimi et al. 2018), decision trees (Herazo-Padilla et al. 2021), and Bayesian approaches (Bellot and van der Schaar 2019)—use structured features as input to model the complex relationships between patient features and clinical outcomes. More recently, deep learning–based frameworks have also been explored to capture these complex dynamics and have shown great potential (Cui et al. 2025b; Arsalan 2025; Yu et al. 2024; Wu, Xiao, and Sun 2023; Zhu et al. 2024). However, deep learning-based approaches remain largely restricted to structured data and rely heavily on statistical correlations between input features and outcome labels. Deep learning-based approaches also often lack interpretability grounded in explicit clinical knowledge and tend to be highly specific to the training dataset, limiting their generalizability to other clinical settings (Jiang et al. 2025, 2024).
NLP and Language Models: Unstructured clinical texts (e.g., physician notes, radiology reports, and nursing progress notes) contain rich information about patients (Johnson et al. 2016; Liu et al. 2024; Jiang et al. 2024). Textual documentation includes clinicians’ reasoning and management plans that are not contained elsewhere in the EHR and which may provide direct and high-value insights. To harness this information, various natural language processing (NLP) methods have been proposed for predictive modeling (Mullenbach et al. 2018; Lipton et al. 2016; Yao, Mao, and Luo 2018). Recently, large language models (LLMs) have demonstrated significant progress across a range of tasks, and researchers have begun exploring their potential for processing unstructured clinical text and supporting clinical reasoning (Liu et al. 2024; Niu et al. 2024; Xu et al. 2024). However, one of the major challenges in applying LLMs to healthcare is the risk of errors and hallucinations (Cui et al. 2025a; Shi et al. 2024). To address this, some works incorporate knowledge graphs to explicitly embed clinical knowledge and mitigate LLM inaccuracies (Soman et al. 2024; Jiang et al. 2024). Despite these advances, integrating both structured EHR data and unstructured clinical text in a unified framework remains a challenging problem (Zhang et al. 2021). Existing approaches—such as prompt engineering (Hu et al. 2024), RAG (Zhu et al. 2024), and data fusion (Cui et al. 2024; Thao et al. 2024)—tend to be ad hoc and lack a systematic method for joint modeling.
| Symbol | Description |
| Structured clinical features | |
| Clinical free-text notes (e.g., physician notes) | |
| Demographic information | |
| Embeddings of similar patient notes | |
| Biomedical knowledge graph, triples | |
| Unstructured text embedding | |
| Knowledge graph community summary embedding | |
| LLM-generated reasoning with context | |
| Augmented LLM training/test inputs | |
| Fine-tuned large language model | |
| LLM-generated textual reasoning, output label | |
| Final machine learning classifier | |
| Concatenated features for | |
| Final binary classification output | |
| LLM fine-tuning loss | |
| Task-specific binary cross-entropy loss |
3 Background
Clinical Datasets. Clinical data used in this study includes both structured and unstructured data collected across patient visits, capturing the patient’s condition over time.
Structured data includes standardized fields such as lab results, vital signs, demographic attributes (e.g., age, sex, ethnicity). These are typically numeric or categorical and readily usable for statistical modeling.
Unstructured data consists of free-text clinical documentation such as physician notes, as well as patient conditions, diagnoses, and prescribed medications.
Additional datasets. PubMed and UMLS (Section 5.1).
Problem statements. To demonstrate the effectiveness of our method, we study two popular clinical tasks: in-hospital mortality prediction and 30-day readmission prediction, e.g., (Jiang et al. 2025; Cai et al. 2016). We define these problems formally after introducing some notation.
Notation: We use to denote a hospital visit by a patient. For each visit , the patient is associated with a set of medical data, , comprising both structured information (e.g., codes, vitals, lab results) and unstructured information (e.g., clinical free-text notes). Our goal is to build a model that predicts a patient’s target status based on their historical visit information, specifically, .
30-day Readmission Problem: The objective is to determine whether the patient is readmitted to the hospital within 30 days following discharge from visit . We define the readmission indicator as:
The goal is to develop a predictive model that estimates using all structured and unstructured data from .
Mortality Prediction Problem. Given the complete set of information for a visit , the objective is to determine the patient’s in-hospital mortality status, denoted as , where:
The goal is to develop a predictive model that accurately estimates based on all available structured and unstructured clinical data associated with visit .
4 KAMELEON Framework
We propose a hybrid framework, KAMELEON, that integrates multimodal EHR data, including structured clinical components and unstructured physician notes (we will provide more details in Section 5.1), and external biomedical knowledge to predict two key clinical outcomes: in-hospital mortality and 30-day readmission. As shown in Figure 1, KAMELEON consists of two components: (1) an unstructured encoder (Section 4.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction, along with its reasoning; and (2) a structured encoder (Section 4.2) that combines multiple time-series corresponding to vitals and tabular datasets (labs, medications, etc), along with the outputs from (i.e., both the prediction and the embedding associated with the reasoning it produces) with static features for downstream prediction. The notations used in Section 4, Algorithm 1, and 2 are explained in Table 1.
4.1 Unstructured Text Encoder,
For each hospital visit , we collect physician-authored clinical notes and extract entities like conditions, procedures, and medications. To enrich context, we use PubMed literature parsed into knowledge triples (subject–relation–object) via an LLM-based extraction pipeline. We retain only triples that appear across patient visits. These triples form a biomedical knowledge graph, serving as an auxiliary source to support LLM reasoning and diagnosis.
Generating Context: The first step of the framework is generating patient context. We use EHR data, including physicians’ notes, patient conditions, prescribed medications, and procedures. Since physician notes can be lengthy and exceed the context window of a small, locally fine-tuned LLM, we summarize them using an LLM. This approach addresses Challenge 2 (Long Context).
Knowledge Graph Retrieval: To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)(Bodenreider 2004), PubMed abstracts, and large language model (LLM)-generated entity-relation triples. UMLS provides standardized biomedical concepts and relationships, and the entity-relation triples are structured facts extracted by the LLM in the form , capturing semantic connections. We apply the Leiden algorithm for community detection (Traag, Waltman, and Eck 2019), which partitions the KG into semantically coherent and well-connected subgraphs. After clustering, we employ a separate LLM to generate a textual summary for each cluster. These summaries are produced by reasoning over the relationships among entities within each cluster, capturing the latent biomedical semantics encoded in the graph. Each summary serves as a high-level abstraction of the biomedical concepts and interactions within a subgraph. We embed these summaries using SentenceTransformer (MiniLM-L6-v2) (Reimers and Gurevych 2019) and retrieve the most relevant ones for each patient by computing semantic similarity with the embedded patient context. This process directly addresses Challenge 3 (Specialized Medical Domains) by enriching patient context with structured, domain-specific knowledge, improving the model’s understanding of specialized medical terminology.
Finding Similar Patients: We provide additional context, by retrieving semantically similar patient visits using precomputed visit-level embeddings and a similarity index using the FAISS (Douze et al. 2024) library with inner-product search on L2-normalized embeddings, which effectively approximates cosine similarity. For each target patient visit, we retrieve the top 50 most similar patients while excluding self-matches and other visits from the same individual. Each retrieved patient is scored by similarity, and we filter them into positive and negative cohorts based on matching or non-matching ground truth labels (e.g., readmission vs. no readmission). The final output includes the top-k positive and negative similar patients (with k=1,2). Unlike KARE (Jiang et al. 2025), we also provide the physician notes of the retrieved similar patients, enabling the language model to leverage more clinical context when assessing patient risk.
Reasoning Module: In this module, we prepare inputs to fine-tune the LLM for clinical prediction. For each patient visit, we create a prompt with the patient’s context with the top-k most similar cases retrieved earlier. These similar cases guide the model by highlighting patterns in clinically comparable scenarios. We also add biomedical knowledge summaries from clustered subgraphs of a PubMed knowledge graph, providing literature-based context. Combining patient data, historical cases, and domain knowledge, we fine-tune the LLM to produce task-specific predictions with interpretable reasoning, supporting each outcome.
Finetuning LLM: We fine-tune a LLaMA-3 8B model using the Unsloth framework (Touvron et al. 2024; Daniel Han and team 2023), which enables memory-efficient training via 4-bit quantization and Low-Rank Adaptation (LoRA) (Hu et al. 2022). Prediction tasks are framed as instruction-following using Alpaca-style prompts with task description, patient context, and optional justification. Each prompt combines clinical notes, retrieved similar cases, and biomedical knowledge summaries. The model is trained via supervised learning to generate both predictions and reasoning. We use limited training steps with gradient accumulation and sequence lengths up to tokens. Unlike KARE, which trains larger models with higher compute, our method uses smaller, quantized models to reduce computational cost while maintaining interpretability and performance. Algorithm 1 outlines the training and inference procedures of . Additionally, an example prompt and its overall structure are provided in the Appendix Section A.3.
4.2 Structured Data Encoder,
We extract structured features from the patient’s visit history during each stay, including:
Time-varying variables: We extract hourly vitals and diagnoses in MIMIC-III, including heart rate, systolic and diastolic blood pressure, mean blood pressure, oxygen saturation, Glasgow Coma Scale (GCS) scores, glucose level, respiratory rate, temperature, weight, and pH.
Static metadata: Demographic and admission features include gender, ethnicity, admission type, location, insurance, language, religion, and hospital outcome.
Diagnoses, procedures, medications: ICD-9 codes, drug names are encoded via one-hot or counts. We compute binary indicators for key comorbidities (e.g., sepsis, infection, organ failure, dementia, cancer, diabetes).
Structured Data Preprocessing: The structured data is first transformed using a discretization step to enforce uniform temporal resolution and impute missing values. This is followed by normalization using precomputed mean and standard deviation statistics over the continuous variables.
Incorporating LLM Output: To augment the structured input, we include the LLM’s prediction and its tokenized reasoning. A Word2Vec (Mikolov et al. 2013) model is trained on this reasoning to generate embeddings. Each paragraph is represented by averaging its word vectors, producing a fixed-size vector. This reasoning vector is concatenated with the output label to form a feature vector for each visit.
Final Integration. The LLM-derived vector is merged with structured input features to create a unified representation, directly addressing Challenge 1 (Multi-modal Information). To reduce dimensionality and suppress noise from high-dimensional embeddings, we apply Principal Component Analysis (PCA) to the combined feature vector.
4.3 Training
We follow a two-stage training procedure. First, we fine-tune the unstructured text encoder using instruction-style prompts built from clinical notes, retrieved similar cases, and external biomedical knowledge. After fine-tuning, we perform Final Integration and use the outputs of as input features to train for final prediction.
In our experiments, we benchmark several ML models for as KAMELEON-X, where X represents logistic regression, random forests, LSTM, LightGBM, MLP, or XGBoost, selected for its effectiveness in capturing clinical patterns. For MLP, we use weighted binary cross-entropy loss
where and are the positive and negative class weights, respectively, used to address class imbalance by giving more emphasis to the minority class. We further use Synthetic Minority Over-sampling Technique (SMOTE) (Chawla et al. 2002) to mitigate class imbalance, addressing Challenge 4 (Highly imbalanced data). The pipeline for is in Algorithm 2.
| Model | Acc | Prec0 | Prec1 | Rec0 | Rec1 | Macro F1 | AUROC | AUPRC | ||
| Task: 30-Day Readmission Prediction | ||||||||||
| LogReg (C=1.0) (Harutyunyan et al. 2019) | ✓ | – | 0.831 | 0.869 | 0.036 | 0.951 | 0.013 | 0.464 | 0.463 | 0.090 |
| MLP (Bopche et al. 2024) | ✓ | – | 0.828 | 0.876 | 0.182 | 0.934 | 0.100 | 0.516 | 0.559 | 0.165 |
| BalancedRF | ✓ | – | 0.760 | 0.970 | 0.070 | 0.780 | 0.430 | 0.490 | 0.673 | 0.066 |
| LSTM (Harutyunyan et al. 2019) | ✓ | – | 0.820 | 0.876 | 0.163 | 0.925 | 0.100 | 0.512 | 0.569 | 0.152 |
| Claude-3.7-Sonnet | – | ✓ | 0.240 | 0.790 | 0.199 | 0.068 | 0.927 | 0.227 | 0.498 | 0.199 |
| Medgemma-4b-it | – | ✓ | 0.350 | 0.770 | 0.190 | 0.270 | 0.690 | 0.350 | 0.480 | 0.190 |
| Llama3-Med42-8B | – | ✓ | 0.390 | 0.800 | 0.210 | 0.360 | 0.670 | 0.410 | 0.510 | 0.210 |
| KARE | – | ✓ | 0.660 | 0.870 | 0.130 | 0.720 | 0.280 | 0.480 | 0.506 | 0.195 |
| KAMELEON-LogReg, (C=1.0) | ✓ | ✓ | 0.833 | 0.869 | 0.037 | 0.953 | 0.013 | 0.519 | 0.130 | 0.148 |
| KAMELEON-LogReg, (C=0.01) | ✓ | ✓ | 0.871 | 0.874 | 0.333 | 0.996 | 0.013 | 0.478 | 0.551 | 0.152 |
| KAMELEON-LSTM | ✓ | ✓ | 0.840 | 0.880 | 0.190 | 0.950 | 0.090 | 0.510 | 0.505 | 0.135 |
| KAMELEON-BalancedRF | ✓ | ✓ | 0.800 | 0.990 | 0.130 | 0.800 | 0.790 | 0.550 | 0.845 | 0.150 |
| KAMELEON-MLP | ✓ | ✓ | 0.910 | 0.970 | 0.160 | 0.940 | 0.280 | 0.580 | 0.820 | 0.138 |
| Task: In-Hospital Mortality Prediction | ||||||||||
| Logistic Regression (Harutyunyan et al. 2019) | ✓ | – | 0.850 | 0.912 | 0.340 | 0.916 | 0.331 | 0.625 | 0.624 | 0.190 |
| LSTM (Harutyunyan et al. 2019) | ✓ | – | 0.690 | 0.800 | 0.260 | 0.800 | 0.250 | 0.530 | 0.560 | 0.240 |
| BalancedRF | ✓ | – | 0.810 | 0.950 | 0.340 | 0.820 | 0.700 | 0.670 | 0.860 | 0.475 |
| LightGBM (Bopche et al. 2024) | ✓ | – | 0.890 | 0.930 | 0.510 | 0.940 | 0.480 | 0.720 | 0.866 | 0.534 |
| MLP (Bopche et al. 2024) | ✓ | – | 0.870 | 0.920 | 0.430 | 0.920 | 0.430 | 0.680 | 0.829 | 0.426 |
| XGBoost | ✓ | – | 0.890 | 0.920 | 0.520 | 0.950 | 0.380 | 0.695 | 0.835 | 0.487 |
| Claude-3.7-Sonnet | – | ✓ | 0.800 | 0.890 | 0.120 | 0.880 | 0.130 | 0.510 | 0.510 | 0.110 |
| Medgemma-4b-it | – | ✓ | 0.120 | 0.950 | 0.100 | 0.020 | 0.990 | 0.120 | 0.510 | 0.110 |
| Llama3-Med42-8B | – | ✓ | 0.160 | 0.950 | 0.120 | 0.100 | 0.970 | 0.190 | 0.530 | 0.120 |
| KARE | – | ✓ | 0.614 | 0.890 | 0.134 | 0.641 | 0.413 | 0.474 | 0.527 | 0.125 |
| KAMELEON-BalancedRF | ✓ | ✓ | 0.880 | 0.930 | 0.490 | 0.934 | 0.492 | 0.710 | 0.876 | 0.543 |
| KAMELEON-LSTM | ✓ | ✓ | 0.730 | 0.820 | 0.430 | 0.840 | 0.390 | 0.620 | 0.740 | 0.350 |
| KAMELEON-lightGBM | ✓ | ✓ | 0.880 | 0.940 | 0.470 | 0.910 | 0.590 | 0.730 | 0.890 | 0.550 |
| KAMELEON-MLP | ✓ | ✓ | 0.900 | 0.940 | 0.550 | 0.940 | 0.550 | 0.750 | 0.890 | 0.600 |
| KAMELEON-XGBoost | ✓ | ✓ | 0.920 | 0.920 | 0.790 | 0.980 | 0.369 | 0.660 | 0.920 | 0.650 |
5 Results
We evaluate our model on the two clinical prediction tasks: in-hospital mortality and 30-day readmission. Our experiments compare performance against strong baselines, including general and medical LLMs, and assess the value of key model components through ablation studies. Some of our results are discussed in Section B Appendix.
5.1 Experimental Setup
Datasets. We use the MIMIC-III dataset (Johnson et al. 2016), which includes structured and unstructured data for over 40,000 ICU patients. It includes structured data (demographics, vitals, labs, admissions, ICD-9-CM codes) and unstructured clinical text (physician notes, discharge summaries, radiology reports). For this study, we focus on physician-authored notes containing clinical reasoning, assessments, and treatment plans. Here, we exclude discharge summaries and notes written after outcomes to prevent label leakage. Only of notes mention hospice, indicating rare explicit terminal indicators.
Biomedical Knowledge. We use abstracts from the annual PubMed Baseline dataset, comprising over 36 million biomedical citation records, to build a medical knowledge graph that enriches LLM input, and reduces hallucinations. We also incorporate UMLS (Bodenreider 2004) to construct concept-centric subgraphs from EHR data.
Dataset Statistics. We include a summary table in the Appendix, showing dataset statistics, indicatin moderate class imbalance for in-hospital mortality (positive) and severe imbalance for 30-day readmission ( positive). To prevent data leakage, train-test splits are patient disjoint, meaning that multiple visits from the same patient do not appear in both sets.
Baselines. We compare against Claude 3.5 Sonnet (Anthropic 2024), MedGemma (Sellergren et al. 2025), LLaMA3-Med (Christophe et al. 2024), and KARE (Jiang et al. 2025), as well as structured-data models including logistic regression, tree-based models, and MLPs used in prior work on MIMIC III (Harutyunyan et al. 2019; Lundberg et al. 2020; Purushotham et al. 2018). All LLMs are evaluated in a zero-shot setting with the same patient-context prompt. KARE uses similar patient retrieval but lacks clinical notes. Our model incorporates retrieved notes for better context. Implementation details are in the Appendix.
Metrics. We report AUROC, AUPRC, accuracy, precision, recall, and macro F1-score. Metrics are computed on a held-out test set with similar standardized preprocessing.
5.2 Performance of KAMELEON
30-Day Readmission Prediction:
Readmission within 30 days is a highly imbalanced task, with only about positive cases in the dataset. This severe imbalance is reflected in the results, where most models achieve high accuracy and precision on the negative class but struggle with recall for the positive (readmitted) class.
As shown in Table 2, our framework with Balanced Random Forest (KAMELEON-BalancedRF) classifier achieves the highest AUROC and notably improves recall on positive cases to , a crucial metric since identifying patients at risk of readmission is clinically imperative. The KAMELEON-MLP model, while achieving the highest overall accuracy and macro F1 , still attains a recall of on positive cases, illustrating the persistent challenge in detecting rare events.
Unstructured LLM-based baselines such as Claude-3.7-Sonnet, MedGemma, LLaMA3-Med, and KARE show substantially lower recall for positives (below 0.3), suggesting that these models struggle to identify the minority class without further fine-tuning or domain-specific adaptation.
To better understand this model’s behavior, we perform SHAP analysis to identify feature importance, and Figure 2 indicates that the model relies primarily on prediction embeddings and lab/vital features for predicting 30-day readmission, highlighting the importance of multimodal inputs.
While prior studies like Morgan et al. (2019) (Morgan et al. 2019) reported AUROCs up to 0.81 for readmission, and general models typically ranged from 0.61 to 0.73 (Dhalluin et al. 2020; Matheny et al. 2021), our multi-modal approach effectively captures complex clinical nuances.
In-Hospital Mortality Prediction: Mortality prediction is less imbalanced, with approximately positive cases. Table 2 reports that KAMELEON-XGBoost and KAMELEON-MLP models achieve high accuracy ( and respectively) and AUROC ( and respectively). Furthermore, for a positive class that, while less imbalanced, is still a minority, the Area Under the Precision-Recall Curve (AUPRC) is a vital metric. Here, the KAMELEON with sets the benchmark, significantly outperforming all other baselines. Unstructured models consistently yield the lowest performance for mortality prediction, AUROC values hover just above random chance (around ), and AUPRCs remain very low (max ), indicating a limited ability to discern between mortality and survival solely from clinical notes. SHAP results (Appendix Figure 8) demonstrate that lab results and vital signs strongly drive predictions, with prediction embeddings playing a less critical role compared to readmission. Overall, Table 2 shows that for both tasks, our multimodal model outperforms all individual structured and unstructured baselines across all metrics.
5.3 Ablation Study
To highlight the impact of reasoning features, we perform an ablation study (Table 8, Appendix). Removing reasoning features alone causes AUROC to drop from to for 30-day readmission. When we further remove demographics and prediction embeddings alongside reasoning, performance falls below , indicating that each component plays a complementary role. Figure 2 visualizes the contribution of reasoning, ’s prediction, and demographics to model performance. We also provide a similar plot for in-hospital mortality prediction in the Appendix. For this task, structured data played a more critical role, with less relative impact from reasoning and unstructured features compared to readmission prediction.
6 Conclusion
In this work, we introduced KAMELEON, a novel framework that effectively integrates multimodal EHR data, including structured clinical features and unstructured physician notes, enhanced by knowledge-augmented LLM reasoning, for robust clinical risk prediction. Our two-stage architecture demonstrated superior performance on both 30-day readmission and in-hospital mortality prediction, achieving a high AUROC score despite severely imbalanced datasets. KAMELEON consistently outperformed existing models, highlighting the synergy achieved by combining these diverse modalities.
This work underscores the significant potential of knowledge-augmented multimodal EHR modeling to enhance early intervention, optimize resource allocation, and improve patient care in complex clinical settings. Future work will focus on further validating KAMELEON’s generalizability across diverse clinical settings and exploring its application to a wider range of predictive healthcare tasks.
Scope and Promise for Social Impact. KAMELEON offers a strong opportunity to reduce avoidable hospital readmissions, a major driver of morbidity, cost, and financial penalties (Fry, Fluck, and Han 2021; Joynt and Jha 2013; Panagiotou et al. 2021; Zuckerman et al. 2017). The greatest impact is expected in high-risk groups, such as Outpatient Parenteral Antimicrobial Therapy (OPAT) patients, who face up to readmission rates (Wolie et al. 2024). OPAT programs are often underfunded and rely on manual, error-prone reviews. By identifying high-risk patients, our model enables focused use of limited resources, improving efficiency and outcomes.
We are exploring deployment in a large regional hospital for such high-impact populations, with later expansion to general readmissions.
Appendix A Additional Details on Dataset
We use the Medical Information Mart for Intensive Care III (MIMIC-III) dataset (Johnson et al. 2016), a large-scale, publicly available critical care database comprising deidentified health-related data associated with over patients who were admitted to the intensive care units (ICUs) of Beth Israel Deaconess Medical Center between and . The dataset was curated to support clinical and epidemiological research and has since become a cornerstone resource in machine learning for healthcare.
MIMIC-III offers a comprehensive view of patient health trajectories by incorporating a diverse range of structured and unstructured data sources. Structured data include demographic information (e.g., age, gender, ethnicity), vital signs recorded at regular intervals (e.g., heart rate, blood pressure, respiratory rate, oxygen saturation), laboratory test results (e.g., glucose, pH, hematocrit), and hospital admission details (e.g., insurance type, admission location, length of stay, discharge disposition). Diagnosis and procedure codes are also available via ICD-9-CM, allowing for condition-based patient cohort extraction.
In addition to structured clinical variables, MIMIC-III contains extensive unstructured clinical text, such as physician notes, radiology reports, discharge summaries, Rehab Services, Echo, ECG , Case Management, General, Social Work, Pharmacy, Consult, and Nursing progress notes. These textual artifacts provide rich narrative context that captures clinical reasoning, impressions, and temporal evolution of patient conditions—information often not fully represented in structured fields.
For this study, we primarily focus on the unstructured physician notes that are authored during and after patient examinations. These notes include clinical assessments, differential diagnoses, treatment plans, and observations made by attending or resident physicians. We extract all available physician notes for each patient and temporally align them using the associated charted timestamps. To ensure semantic fidelity and medical relevance, a subset of the extracted notes is manually reviewed and annotated in collaboration with licensed physicians, enabling accurate downstream processing and interpretation by our model.
A.1 Statistics of Dataset
The datasets for 30-day readmission and in-hospital mortality prediction are highly imbalanced, as only a small fraction of patients who are discharged are readmitted within 30 days. Table 3 shows the percentage of positive cases for both tasks in the training and test datasets.
| Task | Split | #Samples | % Positive |
| In-Hospital Mortality | Train | ||
| Test | |||
| Readmission in 30 days | Train | ||
| Test |
A.2 Features
Table 4 outlines the structured and unstructured features used to train our framework. We incorporate all available patient, and visit-level information, including vitals, lab results, diagnosis and procedure codes, medications, and clinical notes, to provide a comprehensive representation of each patient’s condition. This multimodal input enables the model to learn both high-level clinical states and subtle variations across patients.
Table 5 details the ICD-9 code groupings used to derive comorbidity indicators. These comorbidities are included as structured features, allowing the model to capture critical clinical context such as the presence of chronic diseases (e.g., diabetes, cardiovascular disease) or acute conditions (e.g., sepsis, organ dysfunction). We also found that incorporating diagnosis type and the frequency of related conditions (e.g., number of cancer-related diagnoses) further improved prediction performance, suggesting that leveraging structured clinical coding hierarchies contributes meaningful signal to patient outcome modeling.
| Type | Category | Examples |
| Structured | Static | Age, Gender, Ethnicity, Admission Type, Admission Location, Insurance, Religion |
| Vitals | Heart Rate, Systolic/Diastolic/Mean Blood Pressure, Respiratory Rate, Temperature, Oxygen Saturation, Capillary Refill Rate | |
| Neurological Scores | Glascow Coma Scale: Eye Opening, Motor Response, Verbal Response, Total | |
| Lab Results | pH, Glucose, Fraction Inspired Oxygen | |
| Physical Measures | Height, Weight | |
| Codes | ICD-9 diagnosis codes, Procedure codes (CPT) | |
| Unstructured | Clinical Notes | Discharge summaries, Progress notes, Nursing notes |
| Conditions | respiratory failure, complications of surgical procedures or medical care, aspiration pneumonitis, etc | |
| Procedures | tracheostomy, respiratory, intubation and mechanical ventilation, gastrostomy etc | |
| Medications | hypnotics, sedatives; drugs for constipation, other nutrients in atc,adrenergics, inhalants, etc |
| Category | ICD-9 Prefixes (Examples) |
| Explicit Sepsis | 99591, 99592, 78552, 038 |
| Infection | 001–139 (e.g., 001, 008, 038, 041, 079) |
| Organ Dysfunction | 584, 7855, 570, 572, 51881, 7991, 293 |
| Diabetes | 250 |
| Cardiovascular | 390–429 (e.g., 401, 410, 414, 427) |
| Cancer | 140–189 (e.g., 153, 162, 174, 185) |
| Lung Disease | 490–505 (e.g., 491, 493, 496, 500) |
| Dementia | 290, 294 |
| Kidney Dialysis | 585 |
| Liver Disease | 570, 571, 572 |
| Immune Disorder | 279 |
A.3 Untructured Data Representation for
In the main paper, we described multiple stages of unstructured text encoder, . In this section, we will provide, examples on each step.
Generating Context
In this section, for each patient ’s visit , we utilize information from the MIMIC-III dataset, including ICD codes and physician notes. These are used to extract the patient’s conditions, procedures, and prescribed drugs. Note that not all visits contain physician notes; in such cases, we leave the note field blank. If available, we include the note for the corresponding visit.
For patients with multiple visits, we aggregate information from all prior visits. This means a single entry may contain multiple physician notes concatenated together. However, these notes are often unstructured and may include irrelevant or redundant content, which can lead to context length issues when used as input to large language models (LLMs) during fine-tuning.
To address this, we employ a separate LLM to summarize the concatenated physician notes, producing a concise and clinically relevant summary for each patient history. Figure 3 demonstrates a sample example of a patient context.
Knowledge Graph Retrieval
The first step in our approach involves downloading the latest PubMed dataset and extracting all available abstracts. The goal is to capture general medical knowledge that may support diagnostic reasoning in complex clinical cases.
Concept Generation: After constructing the base context from these abstracts, we treat extracted medical conditions, procedures, and other relevant entities as concepts. In addition to the KARE framework, we further enhance concept extraction from patients’ physician notes by applying a dedicated NLP model to parse physician notes and extract clinical keywords and concepts directly from unstructured text. These are extracted using a biomedical natural language processing pipeline built on scispaCy. Each note is processed with the model to identify entities such as diseases, medications, procedures, symptoms, and conditions. Recognized terms are normalized to lowercase and stored as concepts. For notes where specific biomedical entities are not labeled, relevant noun phrases (longer than three characters) are also included to ensure broader coverage of meaningful medical concepts. These extracted concepts are later combined with structured EHR codes for downstream analysis.
KG Construction: Using these concepts, we perform targeted retrieval from the PubMed abstracts, identifying paragraphs that are semantically aligned with each concept. Once relevant text is retrieved, we employ a large language model (LLM) to extract relational triples (subject, predicate, object), thereby constructing a knowledge graph grounded in biomedical literature. Figure 4 shows a partial snapshot of a knowledge graph constructed from PubMed data, filtered to include only patient-related concepts.
KG structure: This PubMed-derived knowledge graph supplements patient-specific data and improves reasoning by providing broader clinical context. The following triples represent extracted relations involving beta blocking agents from biomedical literature:
(unstable angina pectoris, treated with, beta blocking agents)
beta blocking agents, should be avoided in, non responsive patients)
(beta blocking agents, could be a useful measure in, patients with labile arterial hypertension)
(beta blocking agents, could be a useful measure in, atients with vegetative dysregulation)
(beta blocking agents, could be a useful measure in, patients with hyperkinetic heart syndrome)
(bunitrolol, is a type of, beta blocking agents)
(beta blocking agents, impact, myocardial lactate extraction)
(beta blocking agents, reduce, arterial NEFA levels)
Finding Similar Patients
Here, the model identifies similar patients by leveraging diagnoses, procedures, and medications, along with unstructured data from physician notes. For each patient visit, KARE extracts key clinical concepts and compares them across the cohort to find patients with overlapping medical histories and treatments. This similarity assessment enables more personalized predictions by grounding the model in relevant patient contexts.
| Feature | Patient ID: 12081_1, Visit 0 | Patient ID: 12659_0, Visit 0 |
| Conditions | Congestive heart failure, acute cerebrovascular disease, respiratory failure, COPD, cardiac dysrhythmias, coronary atherosclerosis, conduction disorders | Benign neoplasm, heart valve disorders, hepatitis, hypertension, lipid metabolism disorders |
| Procedures | Pacemaker insertion/revision, blood transfusion, cardiac rhythm conversion | Heart procedures, extracorporeal circulation, blood transfusion |
| Medications | ACE inhibitors, adrenergics (inhalants), corticosteroids, lipid modifiers, cardiac stimulants, potassium supplements, antithrombotics, opioid analgesics | Antiinfectives, ACE inhibitors, beta blockers, calcium channel blockers, diuretics, analgesics, anxiolytics |
| Physician Notes | Continued respiratory failure, CHF, COPD; new lung cancer, heart valve disorders, diagnostic bronchoscopy, new medications | Not available |
| Label | 1 | 0 |
Reasoning Module
A critical component of our framework is the reasoning module, which enables the model to go beyond surface-level pattern matching and perform clinically informed inference. This module integrates the patient data—including summarized clinical notes, and the retrieved biomedical knowledge graph—to generate logical, interpretable explanations supporting the risk predictions.
Specifically, the reasoning module processes extracted clinical concepts and relationships from patient history and external knowledge to construct chains of evidence, mimicking clinician decision-making steps. This approach helps the model distinguish relevant signals from noise in complex and heterogeneous data.
In Figure 6, the reasoning module is illustrated as a dedicated section within the overall prompt structure. It explicitly asks the LLM to produce detailed, medically grounded reasoning before outputting the final risk prediction, which improves both accuracy and interpretability. Ablation studies (Table 8) demonstrate that removing this reasoning component significantly degrades predictive performance, underscoring its essential role.
A.4 Full Example Prompt
Figure 6 illustrates the complete prompt structure used to fine-tune the LLM. The prompt includes (1) a clear task description; (2) the target patient’s electronic health record (EHR) context combining structured features and summarized clinical notes; (3) relevant external biomedical knowledge retrieved from PubMed abstracts represented as a knowledge graph; (4) examples of similar and dissimilar patient cases for comparative context (We show this in Table 6; and (5) reasoning behind the prediction, generated from an LLM.
The prompt concludes with an explicit request for medically grounded reasoning followed by the final risk prediction. This structured input design facilitates interpretable and accurate model outputs.
Sample Inference Output
Table 7 shows a sample output generated by the fine-tuned LLM for the 30-day hospital readmission prediction task. Instead of returning a raw binary value (0 or 1), the model is prompted to generate both a prediction and its reasoning based on the patient’s diagnoses, procedures, medications, and comparisons with similar cases. This structured explanation allows the model to ground its prediction in clinical context, improving reliability and interpretability over naive classification. By incorporating rationale into the output, the LLM demonstrates better alignment with real-world clinical decision-making.
Patient ID: 25070_0
To predict whether Patient ID 25070_0 will be readmitted within 30 days, we analyze the provided EHR context, focusing on conditions, procedures, and medications. 1. Conditions: The patient has multiple serious conditions, including non-Hodgkin’s lymphoma and complications from a device, which are significant risk factors for readmission. However, they also have other conditions that may not be as acute (e.g., asthma, thyroid disorders). 2. Procedures: The patient has undergone several diagnostic procedures, including mastoidectomy, which could lead to complications but are not major surgeries. The patient did not have any major surgical interventions that typically increase readmission risk. 3. Medications: The patient is on a complex regimen including corticosteroids and other systemic medications. The presence of corticosteroids indicates a potential for complications, particularly in the management of their non-Hodgkin’s lymphoma. 4. Similar Patients: The similar patients show that those with chronic obstructive pulmonary disease and heart failure have higher readmission rates. However, Patient ID 25070_0 does not share the same chronic conditions with these patients. 5. Overall Assessment: Although the patient has several chronic and acute conditions, they do not have the same severity or complexity as the patients who have been readmitted. The absence of major recent surgeries and the complexity of their medication regimen suggests a lower risk of readmission. # Prediction # 0 |
Patient ID: 93900_1
To predict whether Patient ID: 93900_1 will be readmitted to the hospital within 30 days, we analyze the patient’s EHR context, focusing on conditions, procedures, and medications. 1. Conditions: The patient has multiple chronic conditions, including congestive heart failure, diabetes mellitus with complications, and various gastrointestinal disorders. These conditions are known to increase the risk of exacerbations and complications, which are significant predictors of readmission. 2. Procedures: The patient has undergone significant procedures, such as debridement and upper gastrointestinal endoscopy. These procedures can lead to complications, necessitating close monitoring and follow-up care, which may contribute to readmission. 3. Medications: The patient is on a complex medication regimen, including beta-blockers and insulins. The complexity of this regimen can lead to adherence issues and potential side effects, which might complicate the patient’s health status and increase the likelihood of readmission. 4. Similar Patients: The similar patients, such as Patient ID: 16994_6, experienced complications from chronic diseases and were readmitted. This indicates a trend in patients with similar conditions and treatment regimens. 5. Overall Risk Assessment: The combination of chronic conditions, significant procedures, and complex medication regimens suggests a high risk of readmission due to potential exacerbations and complications. Given this comprehensive analysis, it is reasonable to conclude that Patient ID: 93900_1 will likely be readmitted to the hospital within 30 days based on the factors outlined above. |
A.5 Structured Data Representation for
We preprocess raw patient records by extracting structured data from MIMIC‑III and organizing it into well-defined prediction datasets. We align each patient to an admission based on predefined cohort criteria, for in‑hospital mortality prediction, we use early admission data, typically features collected within the first 24 hours, to build static representations. These include demographic attributes (such as age, gender, and admission type), physiological measurements (vital signs like heart rate or blood pressure), lab test results, and interventions. All continuous variables are normalized and discretized to create time‑series bins or summary statistics. For readmission prediction, we rely on discharge time data or early records filtered using exclusion criteria (e.g., patients under 18 or with missing data), and construct 30‑day readmission labels by linking consecutive admissions.
To handle missing values, we apply simple imputation strategies to ensure a consistent feature space across all patients. Specifically, missing entries in time-series data are filled using forward-fill and backward-fill methods where possible. For variables still missing after temporal imputation, we use either zero imputation (after normalization) or impute with population-level statistics such as the mean or median, depending on the variable type. This approach ensures that the downstream models receive complete input tensors without discarding any patient records due to sparsity, which is critical in clinical data settings where missingness is common and often informative.
Incorporating Patient Demographics and Admission Details This step involves enriching the dataset with key patient demographics and hospital admission attributes. Then we merge external data tables containing information such as gender, ethnicity, admission type (emergency, elective), insurance, and language. Categorical variables are converted into numeric codes for easier model consumption, while missing values are handled with placeholder categories to maintain data integrity. This results in a structured feature set that captures essential background and administrative details for each patient episode.
Adding Clinical Diagnoses, Procedures, and Medication History Next, the dataset is augmented by integrating clinical history through diagnosis codes, procedure codes, and prescribed medications linked to each hospital admission. For instance, all ICD diagnosis codes recorded during a hospital stay are collected into a list, allowing the creation of binary flags that indicate the presence of major conditions like diabetes or cardiovascular disease. Additionally, common code prefixes are extracted to generate features representing groups of related diagnoses or procedures. Similarly, medication records are summarized into key drug groups. This approach transforms complex clinical histories into a numerical format that models can use to assess patient risk.
Embedding Clinical Reasoning and Merging Predictions The final enrichment involves processing free-text clinical reasoning or explanations associated with predictions. These textual rationales are tokenized and converted into numerical vector embeddings that capture their semantic meaning, for example, representing the sentence “patient shows signs of infection” as a 100-dimensional vector. These embeddings are paired with binary prediction labels and merged back into the main dataset by matching patient episodes. Missing values are filled with neutral defaults to ensure consistent data structure. The combined result is a dataset where each patient episode has both predictive labels and rich, encoded reasoning features to support improved downstream modeling or interpretation.
Appendix B Additional results
Ablation on Readmission Table 8 presents an ablation study highlighting the performance impact of removing specific input features and components from models and for the readmission prediction task. The variations include removal of reasoning components, demonstrations, and predicted features.
KAMELEON achieves an accuracy of 0.80 and a Macro F1 score of 0.55, indicating a balanced performance across both classes. When the reasoning component is removed, we observe a spike in accuracy (0.92), but a sharp drop in recall for the positive class (Rec1 = 0.13), showing that the model becomes overly biased toward the negative class. This is further reflected in the reduced AUPRC and AUROC values.
Removing reasoning, demonstrations, and predicted features together leads to the worst performance (Macro F1 = 0.39), confirming that these components are critical for model generalization, especially in correctly identifying positive cases. The stepwise removal of features helps isolate the contribution of each component, underlining the importance of structured reasoning and prior knowledge in our approach.
Ablation on Mortality However, the SHAP values (Figure 8) show that the prediction embedding has a smaller impact compared to lab vitals, which contrasts with the readmission task. Additionally, we observe a higher AUROC for in-hospital mortality even without reasoning, exceeding . This high AUROC score suggests that predicting mortality is comparatively easier, likely due to the dataset being less imbalanced. Given access to lab vitals and diagnostic information, the model can more easily learn patterns associated with terminal illness.
| Model | Acc | Prec0 | Prec1 | Rec0 | Rec1 | Macro F1 | AUROC | AUPRC |
| our setup | 0.80 | 0.99 | 0.13 | 0.80 | 0.77 | 0.55 | 0.844 | 0.147 |
| without reasoning | 0.92 | 0.97 | 0.11 | 0.96 | 0.13 | 0.54 | 0.687 | 0.082 |
| without demo, reasoning, pred | 0.54 | 0.97 | 0.05 | 0.54 | 0.58 | 0.39 | 0.559 | 0.0433 |
| without reasoning, pred | 0.57 | 0.97 | 0.05 | 0.57 | 0.62 | 0.41 | 0.594 | 0.048 |
Analysis of distribution (In-Hospital Mortality)
Figure 9 shows predicted probabilities for in-hospital mortality within 30 days, separated by true class: survivors (blue) and deceased (red). The dashed line marks the optimal threshold (0.26) balancing sensitivity and specificity.
Most survivors cluster near zero probability, reflecting strong model confidence, while deceased cases spread across a wider range, showing prediction uncertainty. Overlap between classes causes some misclassifications, highlighting the challenge of predicting this rare event. Despite this, the clear separation and tight clustering of survivors demonstrate the model’s strong ability to distinguish between classes, supporting the usefulness of the selected threshold.
Class Separability Analysis Using t-SNE (In-hospital Mortality Task)
Figure 10 shows a two-dimensional t-SNE embedding of the test data, where the samples are colored by their true class labels (Class 0 in blue and Class 1 in red). The visualization reveals that the majority of the data forms a consistent, structured manifold dominated by Class 0 points, with only a small fraction of Class 1 points distributed across the embedding. Notably, a compact cluster of Class 1 samples appears on the right-hand side, indicating localized patterns that can be exploited by advanced models. This structure suggests that although Class 1 is relatively sparse, it exhibits distinct feature signatures in specific regions, which the proposed KAMELEON‑X model is designed to capture effectively, contributing to improved predictive performance.
Using Anomaly Detection Method for Readmission in 30 days: Given the highly imbalanced nature of the 30-day readmission data, where only approximately of hospital visits result in readmission, we explored whether this task could be framed as an anomaly detection problem. Specifically, we investigated whether the minority class (readmitted patients) exhibits anomalous patterns relative to the majority class.
To evaluate this, we applied unsupervised anomaly detection methods, such as the Isolation Forest algorithm, to identify readmissions as statistical outliers based on the available clinical features. Table 9 shows the distribution of positive and negative samples in the training and test sets, along with the number of anomalies detected by the model.
Despite the appeal of this approach, anomaly detection methods did not perform well for this task. The results indicate that the readmitted patients do not form distinct statistical outliers in the feature space, making it difficult for unsupervised models to reliably separate them from the majority class. This suggests that the features available do not naturally differentiate readmissions as anomalous events, and that supervised learning approaches are more appropriate for this problem.
| Positive | Negative | Anomalies Detected | Actual positive | |
| Train Set | 718 | 17185 | 717 | 33 |
| Test Set | 123 | 3113 | 130 | 3 |

### v2

@@
-Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling
+Improving Hospital Risk Prediction with Knowledge-Augmented Multimodal EHR Modeling
 Abstract
@@
 1 Introduction
-Appropriate use of clinical prediction tools for early identification of high-risk patients for different conditions allows for clinical decision making, timely interventions, escalation of care, intensive monitoring, and identification of gaps in outpatient management.
-e.g., (Cai et al.
-2016; Kong, Lin, and Hu 2020; Brajer et al.
-2020; Kansagara et al.
-2011; Mahmoudi et al.
-2020).
-For instance, readmission within a short period is a priority under many regulatory frameworks and value-based care models, where high readmission rates may lead to financial penalties; e.g., (Upadhyay, Stephenson, and Smith 2019; Clement et al.
-2017).
-Therefore, good models of short-term risk prediction (e.g., 30 days) can guide targeted interventions such as more detailed discharge instructions, closer post-discharge monitoring, or referrals to transitional care programs.
-While traditional prediction tools relied on logistic regression for risk assessment, machine learning methods are increasingly applied to clinical prediction tasks.
-There has been a lot of work on developing diverse kinds of machine learning methods for such risk prediction questions using Electronic Health Record (EHR) data, which contain rich information on patient health, e.g., (Cai et al.
-2016; Jiang et al.
-2025; Qiu et al.
-2025; Yang et al.
-2023; Brisimi et al.
-2018; Bellot and van der Schaar 2019; Cui et al.
-2025b; Arsalan 2025; Yu et al.
-2024; Wu, Xiao, and Sun 2023; Zhu et al.
-2024).
-Most of this work focuses on structured EHR data, which spans various types and representations (Hartvigsen et al.
-2018), including admission/discharge information, procedures and interventions, medications, lab orders and results, billing codes (e.g., ICD, CPT), physiological time-series (e.g., vital signs), imaging data, and clinical documentation such as progress notes or discharge summaries.
-When used for machine learning approaches to clinical prediction, unstructured data from clinical notes has been used in a fairly simple manner, such as bag of words, TF-IDF (Wu and Zhang 2021; Huang, Osorio, and Sy 2019) to facilitate the use of conventional machine learning methods for clinical tasks.
+Appropriate use of clinical prediction tools for early identification of high-risk patients for different conditions allows for clinical decision making, timely interventions, escalation of care, intensive monitoring, and identification of gaps in outpatient management, e.g., [1, 2, 3, 4, 5].
+For instance, readmission within a short period is a priority under many regulatory frameworks and value-based care models, where high readmission rates may lead to financial penalties, e.g., [6, 7].
+Therefore, effective models of short-term risk prediction (e.g., 30 days) can guide targeted interventions, including more detailed discharge instructions, closer post-discharge monitoring, or referrals to transitional care programs.
+While traditional prediction tools relied on simple statistical models, e.g., regression and decision trees [8, 9], for risk assessment, more complex machine learning methods are increasingly applied to clinical prediction tasks, e.g., [10, 11, 12, 13, 14].
+In this work, we focus on two commonly studied clinical problems: (1) 30-day readmission prediction, which determines whether a patient will be readmitted to the hospital within 30 days after discharge, and (2) mortality prediction, which determines the patient’s in-hospital mortality status.
+There has been a lot of work on developing diverse kinds of machine learning methods for these problems using Electronic Health Record (EHR) data, which contain rich information on patient health, e.g., [1, 15, 16, 17, 8, 18, 10, 11, 12, 13, 14].
+Most of this work focuses on structured EHR data, which spans various types and representations [19], including admission/discharge information, procedures and interventions, medications, lab orders and results, billing codes (e.g., ICD, CPT), physiological time-series (e.g., vital signs), imaging data, and clinical documentation such as progress notes or discharge summaries.
+Unstructured data from clinical notes has been used in a fairly simple manner, such as bag of words, TF-IDF [20, 21] to facilitate the use of conventional machine learning methods for clinical tasks.
 Clinical notes are complex and poorly structured, which limits their utility in clinical informatics tasks, even when using advanced natural language processing (NLP) techniques.
-While large language models (LLMs) offer a powerful means to process such notes, especially when combined with large biomedical datasets to capture richer semantics beyond keywords and embeddings (Shi et al.
-2024; Li et al.
-2024; Singhal et al.
-2023; Jiang et al.
-2025), they still face significant limitations.
-Hallucinations, factual inaccuracies, and a lack of domain grounding make their use in clinical tasks challenging.
-Even standard methods such as retrieval-augmented generation (RAG) have produced inaccurate predictions in clinical settings (Shi et al.
-2024; Jiang et al.
-2025).
-For instance, models such as Med-PaLM (Singhal et al.
-2023) and EHR-Chain-of-Thought (Jiang et al.
-2025) exhibit strong language generation capabilities but frequently misinterpret similar-sounding medical terms.
-Similarly, RAG-based systems like EHRAgent (Shi et al.
-2024) struggle when the retrieved context fails to align with a patient’s history.
-Recent approaches have attempted to enhance LLMs with structured knowledge via graph-based retrieval (e.g., GraphRAG), but their performance remains limited due to a lack of explicit reasoning.
-Efforts such as MedRetriever (Ye et al.
-2021), GraphCare (Jiang et al.
-2024), RAM-EHR (Xu et al.
-2024), EHR-KnowGen (Niu et al.
-2024), and EMERGE (Zhu et al.
-2024) inject knowledge into patient representations, yet still fall short of leveraging the full reasoning potential of LLMs.
+While large language models (LLMs) offer a powerful means to process such notes, especially when combined with large biomedical datasets to capture richer semantics beyond keywords and embeddings [22, 23, 24, 15], they still face significant limitations, such as hallucinations, factual inaccuracies, and inadequate domain grounding [22, 15].
+For instance, models such as Med-PaLM [24] exhibit strong language generation capabilities but frequently misinterpret similar-sounding medical terms.
+Recent approaches have attempted to enhance LLMs with structured knowledge via graph-based retrieval (e.g., GraphRAG), but their performance remains limited due to a lack of explicit reasoning, e.g., [25, 26, 27, 28, 14].
 In recent work, Jiang et al.
-(Jiang et al.
-2025) developed KARE, a graphRAG and context augmentation approach, for clinical prediction tasks, which address many challenges associated with using LLMs for clinical tasks on the MIMIC dataset.
-However, the performance of all prior methods remains limited, and these methods are not usable in clinical decision making.
-Challenges.
-Clinical tasks using EHRs present the following non-trivial challenges: (1) Multi-modal Information: Clinical records in EHRs include both structured and unstructured data.
-How can we effectively handle both modalities within a single processing pipeline?
-(2) Long Context: Clinical documents are typically verbose and repetitive.
-How can we extract relevant information efficiently while avoiding distractions from redundant content?
-(3) Specialized Medical Domains: EHRs involve domain-specific terminology, such as drugs and treatments.
-How can we ensure accurate interpretation of such specialized medical information?
-(4) Highly imbalanced data: Clinical datasets are typically very imbalanced, e.g., about 4% positive for 30-day readmission and about 13% positive for mortality, which has a severe impact on the performance of most ML methods.
-We will address these questions in the framework Section 4.
-Our contributions.
-1. We develop a novel framework, KAMELEON (Knowledge-Augmented Multimodal EHR LEarning for Outcome predictioN)
-that integrates multimodal EHR data (including structured clinical components and unstructured physician notes), and external biomedical knowledge.
-KAMELEON consists of two components (Figure 1):
-(1) a graphRAG and knowledge-augmented reasoning approach, which builds on (Jiang et al. 2025), and outputs a prediction and reasoning; and
-(2) a structured machine learning module which uses tabular and time-series datasets in the EHR, along with the outputs of the LLM based component.
+[15] developed KARE, a graphRAG and context augmentation approach, for clinical prediction tasks, which address many challenges associated with using LLMs for clinical tasks on the MIMIC dataset.
+However, the performance of all prior methods remains limited because clinical tasks using EHRs present many non-trivial challenges: (1) Multi-modality of clinical data: The presence of both structured data and unstructured text requires methods capable of effectively learning from both modalities.
+(2) Long-context textual data: Clinical text often contains a mix of specialized medical terminology and informal or colloquial expressions, making information retrieval challenging.
+(3) Severe class imbalance: Prediction tasks are typically highly imbalanced.
+For example, only about 4% of patients are readmitted within 30 days, resulting in heavily skewed training data.
+Here, we develop a novel framework, KAMELEON (Knowledge-Augmented Multimodal EHR LEarning for Outcome predictioN)
+that addresses the limitations of prior work by integrating multimodal EHR data (including structured clinical components and unstructured physician notes), and external biomedical knowledge.
+KAMELEON consists of two components (as shown in Figure 1):
+(1) An unstructured model (Section 4.2.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction for a patient, along with its reasoning; this extends the approach of [15].
+Physician notes in EHRs can be lengthy and exceed the context window, and these are summarized using an LLM, and used as context.
+To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)[29], PubMed abstracts, and large language model (LLM)-generated entity-relation triples.
+KG is partitioned into semantically coherent and well-connected clusters, and textual summary generated by an LLM for the most relevant clusters for each patient cluster, are used to enrich the context.
+Furthermore, labeled context is added by identifying semantically similar patient visits, which is used to fine-tune the LLM.
+Finally, produces a prediction for the patient, along with a reasoning.
+(2) A structured model (Section 4.2.2), which extracts structured features from the patient’s EHR for the stay, including
+(a) Static demographic and admission data,
+(b) time-varying vitals (which are normalized and summarized, when used as features), and
+(c) diagnoses, procedures and medications.
+In addition, includes the LLM’s prediction and its tokenized reasoning transformed into an embedding, as inputs.
+Finally, different kinds of standard machine learning methods are trained in using these inputs.
+We first train separately, and use the LLM outputs to train .
+We demonstrate the effectiveness of KAMELEON for the 30-day readmission risk and mortality prediction tasks, which have been studied extensively, both using MIMIC-III and other private EHRs from specific hospitals.
+We compare the performance with a number of structured ML and LLM baselines, with respect to multiple metrics (discussed in Section 2).
+KAMELEON consistently outperforms all prior work on MIMIC-III datasets.
+It also shows clear gains over the strongest unstructured LLM baseline (Llama3-Med42-8B).
+The only other prior work which has similar performance for 30-day readmission [30] is on a Norwegian EHR dataset, which is significantly less imbalanced ( readmission positive rate, instead of in the case of MIMIC-III).
+In summary, KAMELEON is the first systematic framework to enhance the power of LLMs for healthcare prediction tasks through graph-guided knowledge retrieval combined with structured machine learning methods.
+We expect this framework to be readily applicable to other clinical questions beyond those examined in this study.
+| Model | Acc | NPV | Precision (PPV) | Specificity | Sensitivity | Macro F1 | AUROC | AUPRC | ||
+| Task: 30-Day Readmission Prediction | ||||||||||
+| LogReg (C=1.0) [31] | ✓ | – | 0.831 | 0.869 | 0.036 | 0.951 | 0.013 | 0.464 | 0.463 | 0.090 |
+| MLP [30] | ✓ | – | 0.828 | 0.876 | 0.182 | 0.934 | 0.100 | 0.516 | 0.559 | 0.165 |
+| BalancedRF | ✓ | – | 0.760 | 0.970 | 0.070 | 0.780 | 0.430 | 0.490 | 0.673 | 0.066 |
+| LSTM [31] | ✓ | – | 0.820 | 0.876 | 0.163 | 0.925 | 0.100 | 0.512 | 0.569 | 0.152 |
+| Claude-3.7-Sonnet† [32] | – | ✓ | 0.240 | 0.790 | 0.199 | 0.068 | 0.927 | 0.227 | 0.498 | 0.199 |
+| Medgemma-4b-it† [33] | – | ✓ | 0.350 | 0.770 | 0.190 | 0.270 | 0.690 | 0.350 | 0.480 | 0.190 |
+| Llama3-Med42-8B† [34] | – | ✓ | 0.390 | 0.800 | 0.210 | 0.360 | 0.670 | 0.410 | 0.510 | 0.210 |
+| – | ✓ | 0.660 | 0.870 | 0.130 | 0.720 | 0.280 | 0.480 | 0.506 | 0.195 | |
+| KARE [15] | – | ✓ | 0.271 | 0.785 | 0.191 | 0.131 | 0.851 | 0.269 | 0.491 | 0.191 |
+| KAMELEON-LogReg, (C=1.0) | ✓ | ✓ | 0.833 | 0.869 | 0.037 | 0.953 | 0.013 | 0.519 | 0.130 | 0.148 |
+| KAMELEON-LogReg, (C=0.01) | ✓ | ✓ | 0.871 | 0.874 | 0.333 | 0.996 | 0.013 | 0.478 | 0.551 | 0.152 |
+| KAMELEON-LSTM | ✓ | ✓ | 0.840 | 0.880 | 0.190 | 0.950 | 0.090 | 0.510 | 0.505 | 0.135 |
+| KAMELEON-BalancedRF | ✓ | ✓ | 0.800 | 0.990 | 0.130 | 0.800 | 0.790 | 0.550 | 0.845 | 0.150 |
+| KAMELEON-MLP | ✓ | ✓ | 0.910 | 0.970 | 0.160 | 0.940 | 0.280 | 0.580 | 0.820 | 0.138 |
+| Task: In-Hospital Mortality Prediction | ||||||||||
+| Logistic Regression [31] | ✓ | – | 0.850 | 0.912 | 0.340 | 0.916 | 0.331 | 0.625 | 0.624 | 0.190 |
+| LSTM [31] | ✓ | – | 0.690 | 0.800 | 0.260 | 0.800 | 0.250 | 0.530 | 0.560 | 0.240 |
+| BRF | ✓ | – | 0.810 | 0.950 | 0.340 | 0.820 | 0.700 | 0.670 | 0.860 | 0.475 |
+| LightGBM [30] | ✓ | – | 0.890 | 0.930 | 0.510 | 0.940 | 0.480 | 0.720 | 0.866 | 0.534 |
+| MLP [30] | ✓ | – | 0.870 | 0.920 | 0.430 | 0.920 | 0.430 | 0.680 | 0.829 | 0.426 |
+| XGBoost | ✓ | – | 0.890 | 0.920 | 0.520 | 0.950 | 0.380 | 0.695 | 0.835 | 0.487 |
+| Claude-3.7-Sonnet† [32] | – | ✓ | 0.800 | 0.890 | 0.120 | 0.880 | 0.130 | 0.510 | 0.510 | 0.110 |
+| Medgemma-4b-it† [33] | – | ✓ | 0.120 | 0.950 | 0.100 | 0.020 | 0.990 | 0.120 | 0.510 | 0.110 |
+| Llama3-Med42-8B† [34] | – | ✓ | 0.160 | 0.950 | 0.120 | 0.100 | 0.970 | 0.190 | 0.530 | 0.120 |
+| – | ✓ | 0.614 | 0.890 | 0.134 | 0.641 | 0.413 | 0.474 | 0.527 | 0.125 | |
+| KARE [15] | – | ✓ | 0.639 | 0.885 | 0.129 | 0.678 | 0.353 | 0.478 | 0.515 | 0.122 |
+| KAMELEON-BRF | ✓ | ✓ | 0.880 | 0.930 | 0.490 | 0.934 | 0.492 | 0.710 | 0.876 | 0.543 |
+| KAMELEON-LSTM | ✓ | ✓ | 0.730 | 0.820 | 0.430 | 0.840 | 0.390 | 0.620 | 0.740 | 0.350 |
+| KAMELEON-lightGBM | ✓ | ✓ | 0.880 | 0.940 | 0.470 | 0.910 | 0.590 | 0.730 | 0.890 | 0.550 |
+| KAMELEON-MLP | ✓ | ✓ | 0.900 | 0.940 | 0.550 | 0.940 | 0.550 | 0.750 | 0.890 | 0.600 |
+| KAMELEON-XGBoost | ✓ | ✓ | 0.920 | 0.920 | 0.790 | 0.980 | 0.369 | 0.660 | 0.920 | 0.650 |
+2 Results
+We evaluate our model on the two clinical prediction tasks: in-hospital mortality and 30-day readmission.
+Our experiments compare performance against strong baselines, including general and medical LLMs, and assess the value of key model components through ablation studies.
+2.1 Experimental Setup
+Datasets.
+We use the MIMIC-III dataset [35], which includes structured and unstructured data for over 40,000 ICU patients.
+It includes structured data (demographics, vitals, labs, admissions, ICD-9-CM codes) and unstructured clinical text (physician notes, discharge summaries, radiology reports).
+For this study, we focus on physician-authored notes containing clinical reasoning, assessments, and treatment plans.
+Here, we exclude discharge summaries and notes written after outcomes to prevent label leakage.
+Only of notes mention hospice, indicating rare explicit terminal indicators.
+However, a limitation of MIMIC-III for the readmission task is the inability to distinguish planned or elective readmissions and interfacility transfers, which may inflate the count of “avoidable” readmissions.
+Biomedical Knowledge.
+We use abstracts from the annual PubMed Baseline dataset, comprising over 36 million biomedical citation records, to build a medical knowledge graph that enriches LLM input, and reduces hallucinations.
+We also incorporate UMLS [29] to construct concept-centric subgraphs from EHR data.
+Dataset Statistics.
+We include a summary table in the Appendix, showing dataset statistics, indicatin moderate class imbalance for in-hospital mortality ( positive) and severe imbalance for 30-day readmission ( positive).
+To prevent data leakage, train-test splits are patient disjoint, meaning that multiple visits from the same patient do not appear in both sets.
+Baselines.
+We compare against Claude 3.5 Sonnet [32], MedGemma [33], LLaMA3-Med [34], and KARE [15], as well as structured-data models including logistic regression, tree-based models, and MLPs used in prior work on MIMIC III [31, 36, 37].
+All LLMs are evaluated in a zero-shot setting with the same patient-context prompt.
+KARE uses similar patient retrieval but lacks clinical notes.
+Our model incorporates retrieved notes for better context.
+Implementation details are in the Appendix.
+Metrics.
+We evaluate model performance using a comprehensive set of measures.
+AUROC and AUPRC capture discrimination ability across thresholds, while overall accuracy reflects the proportion of correct predictions.
+For class-specific evaluation, we report precision (positive predictive value, PPV), recall (sensitivity), negative predictive value (NPV), and specificity.
+These capture both the model’s ability to correctly identify positive cases (sensitivity, PPV) and its reliability on negatives (specificity, NPV).
+Finally, macro F1-score balances precision and recall across classes, ensuring that performance on the minority positive class is not overshadowed by the majority class.
+All metrics are computed on a held-out test set using standardized preprocessing to ensure comparability across models.
+2.2 Performance of KAMELEON
+As we will later discuss in Section 4, we explore different kinds of standard ML methods in for making the final prediction, using all the integrated inputs.
+We refer to the corresponding algorithm as KAMELEON-X, where X represents Balanced-RF (balanced random forest), logistic regression, random forests, LSTM, LightGBM, MLP, or XGBoost.
+2.2.1 30-Day Readmission Prediction
+Readmission within 30 days is a highly imbalanced task, with only about positive cases in the dataset.
+This severe imbalance is reflected in the results, where most models achieve high accuracy and precision on the negative class but struggle with recall for the positive (readmitted) class.
+As shown in Table 1, our framework with Balanced Random Forest (KAMELEON-BalancedRF) classifier achieves the highest AUROC and notably improves recall on positive cases to , a crucial metric since identifying patients at risk of readmission is clinically imperative.
+The KAMELEON-MLP model, while achieving the highest overall accuracy and macro F1 , still attains a sensitivity of on positive cases, illustrating the persistent challenge in detecting rare events.
+Unstructured LLM-based baselines such as Claude-3.7-Sonnet, MedGemma, LLaMA3-Med, and KARE show substantially lower sensitivity for positives (below 0.3), suggesting that these models struggle to identify the minority class without further fine-tuning or domain-specific adaptation.
+In Table 1, KAMELEON-BalancedRF achieves a precision of 0.13, meaning that about one in eight patients flagged as high-risk were actually readmitted.
+Recall (sensitivity) is 0.79, indicating that the model correctly identifies nearly 8 out of 10 true readmissions—a clinically critical result.
+The F1-score of 0.55 reflects the balance between precision and recall.
+The Negative Predictive Value (NPV) is 0.99, showing that almost all patients predicted as low-risk were indeed not readmitted.
+Specificity is 0.80, meaning the model correctly classifies 8 in 10 patients who were not readmitted as low-risk.
+To better understand this model’s behavior, we perform SHAP analysis to identify feature importance, and Figure 2 indicates that the model relies primarily on prediction embeddings and lab/vital features for predicting 30-day readmission, highlighting the importance of multimodal inputs.
+While prior studies like Morgan et al.
+(2019) [38] reported AUROCs up to 0.81 for readmission, and general models typically ranged from 0.61 to 0.73 [39, 40], our multi-modal approach effectively captures complex clinical nuances.
+Relative importance of different classes of inputs.
+We conduct an ablation study for the readmission task, where we retrain KAMELEON after dropping different components— the retraining and prediction models from and the demographics used in (Figure 2 and Table 2).
+We find that the reasoning component output by is very significant, and affects multiple metrics beyond AUROC.
+In the full model, KAMELEON achieves balanced performance (Acc = 0.80, Macro-F1 = 0.55, AUROC = 0.844, AUPRC = 0.147) with both high specificity (0.80) and sensitivity (0.77).
+When we drop the reasoning component, sensitivity falls by over 80% (falling to 0.13), and AUPRC is nearly halved, revealing strong bias toward the majority class; however, there is a gain in accuracy (rising to 0.92).
+Removing the reasoning component from KAMELEON drops performance from 0.844 to 0.7, a 17% decline in AUROC, highlighting the critical role of the finetuned LLM’s reasoning in risk prediction.
+Removing reasoning and prediction (output from ) causes intermediate degradation, with AUROC falling by , consistent with the loss of calibrated probability signals and semantic rationale.
+Eliminating reasoning together with demographics and prediction yields the sharpest overall decline, with Macro-F1 dropping by and AUROC by more than (to 0.663), confirming the complementary value of these components.
+These analyses collectively demonstrate that reasoning substantially improves minority-class detection: it boosts sensitivity and AUPRC by more than threefold compared to variants without it, while also preventing misleading accuracy gains driven solely by the dominant negative class.
+Figure 2 visualizes these contributions, showing that reasoning, ’s prediction embedding, and demographics each add critical and non-redundant signals for accurate readmission risk estimation.
+| Model | Acc | NPV | Precision (PPV) | Specificity | Sensitivity | Macro F1 | AUROC | AUPRC |
+| KAMELEON | 0.80 | 0.99 | 0.13 | 0.80 | 0.77 | 0.55 | 0.844 | 0.147 |
+| without reasoning | 0.94 | 0.96 | 0.07 | 0.97 | 0.06 | 0.52 | 0.699 | 0.078 |
+| without demographic, reasoning, prediction | 0.73 | 0.97 | 0.06 | 0.74 | 0.45 | 0.48 | 0.663 | 0.061 |
+| without reasoning, prediction | 0.74 | 0.97 | 0.09 | 0.75 | 0.55 | 0.50 | 0.691 | 0.062 |
+Sample Inference Output from .
+Table 3 shows a sample output generated by the fine-tuned LLM for the 30-day hospital readmission prediction task.
+Instead of returning a raw binary value (0 or 1), the model is prompted to generate both a prediction and its reasoning based on the patient’s diagnoses, procedures, medications, and comparisons with similar cases.
+This structured explanation allows the model to ground its prediction in a clinical context, improving reliability and interpretability over naive classification.
+By incorporating rationale into the output, the LLM demonstrates better alignment with real-world clinical decision-making.
+Patient ID: 25070_0
+To predict whether Patient ID 25070_0 will be readmitted within 30 days, we analyze the provided EHR context, focusing on conditions, procedures, and medications.
+1.
+Conditions: The patient has multiple serious conditions, including non-Hodgkin’s lymphoma and complications from a device, which are significant risk factors for readmission.
+However, they also have other conditions that may not be as acute (e.g., asthma, thyroid disorders).
 2.
-We demonstrate the effectiveness of KAMELEON for two clinical tasks: 30-day readmission risk and mortality prediction, which have been studied extensively, both using MIMIC-III and other private EHRs from specific hospitals.
-We compare the performance with a number of structured ML and LLM baselines, with respect to multiple metrics.
-KAMELEON achieves significantly better performance for these tasks than all prior work on MIMIC-III datasets.
-We observe a relative improvement in the AUROC of approximately for the 30-day clinical readmission prediction task, and around for the mortality prediction task.
-Similarly, we achieve a relative AUROC improvement of for 30-day readmission and for in-hospital mortality over the best unstructured LLM baseline (Llama3-Med42-8B).
-Our work outperforms clinical score methods as well.
-The only other prior work which has similar performance for 30-day readmission (Bopche et al.
-2024) is on a Norwegian EHR dataset, which is significantly less imbalanced ( readmission positive rate, instead of in the case of MIMIC-III).
-In summary, KAMELEON is the first systematic framework to enhance the power of LLMs for healthcare prediction tasks by using graph-guided knowledge retrieval with structured machine learning methods; we expect this framework to be easily applicable to other clinical questions beyond the ones we study here.
-Note: Many details of the methodology and additional results are included in the Appendix for brevity.
-2 Related Works
-Clinical Prediction using Statistical and Machine Learning Methods: Structured data collected from electronic health records (EHRs) contain valuable information about patients and can be instrumental in capturing patient states and predicting medical outcomes (e.g., mortality) (Jiang et al.
-2025; Qiu et al.
-2025; Yang et al.
-2023).
-To leverage these data, traditional machine learning models—such as linear regression (Brisimi et al.
-2018), decision trees (Herazo-Padilla et al.
-2021), and Bayesian approaches (Bellot and van der Schaar 2019)—use structured features as input to model the complex relationships between patient features and clinical outcomes.
-More recently, deep learning–based frameworks have also been explored to capture these complex dynamics and have shown great potential (Cui et al.
-2025b; Arsalan 2025; Yu et al.
-2024; Wu, Xiao, and Sun 2023; Zhu et al.
-2024).
-However, deep learning-based approaches remain largely restricted to structured data and rely heavily on statistical correlations between input features and outcome labels.
-Deep learning-based approaches also often lack interpretability grounded in explicit clinical knowledge and tend to be highly specific to the training dataset, limiting their generalizability to other clinical settings (Jiang et al.
-2025, 2024).
-NLP and Language Models: Unstructured clinical texts (e.g., physician notes, radiology reports, and nursing progress notes) contain rich information about patients (Johnson et al.
-2016; Liu et al.
-2024; Jiang et al.
-2024).
-Textual documentation includes clinicians’ reasoning and management plans that are not contained elsewhere in the EHR and which may provide direct and high-value insights.
-To harness this information, various natural language processing (NLP) methods have been proposed for predictive modeling (Mullenbach et al.
-2018; Lipton et al.
-2016; Yao, Mao, and Luo 2018).
-Recently, large language models (LLMs) have demonstrated significant progress across a range of tasks, and researchers have begun exploring their potential for processing unstructured clinical text and supporting clinical reasoning (Liu et al.
-2024; Niu et al.
-2024; Xu et al.
-2024).
-However, one of the major challenges in applying LLMs to healthcare is the risk of errors and hallucinations (Cui et al.
-2025a; Shi et al.
-2024).
-To address this, some works incorporate knowledge graphs to explicitly embed clinical knowledge and mitigate LLM inaccuracies (Soman et al.
-2024; Jiang et al.
-2024).
-Despite these advances, integrating both structured EHR data and unstructured clinical text in a unified framework remains a challenging problem (Zhang et al.
-2021).
-Existing approaches—such as prompt engineering (Hu et al.
-2024), RAG (Zhu et al.
-2024), and data fusion (Cui et al.
-2024; Thao et al.
-2024)—tend to be ad hoc and lack a systematic method for joint modeling.
+Procedures: The patient has undergone several diagnostic procedures, including mastoidectomy, which could lead to complications but are not major surgeries.
+The patient did not have any major surgical interventions that typically increase readmission risk.
+3.
+Medications: The patient is on a complex regimen including corticosteroids and other systemic medications.
+The presence of corticosteroids indicates a potential for complications, particularly in the management of their non-Hodgkin’s lymphoma.
+4.
+Similar Patients: The similar patients show that those with chronic obstructive pulmonary disease and heart failure have higher readmission rates.
+However, Patient ID 25070_0 does not share the same chronic conditions with these patients.
+5.
+Overall Assessment: Although the patient has several chronic and acute conditions, they do not have the same severity or complexity as the patients who have been readmitted.
+The absence of major recent surgeries and the complexity of their medication regimen suggests a lower risk of readmission.
+# Prediction # 0 |
+Patient ID: 93900_1
+To predict whether Patient ID: 93900_1 will be readmitted to the hospital within 30 days, we analyze the patient’s EHR context, focusing on conditions, procedures, and medications.
+1.
+Conditions: The patient has multiple chronic conditions, including congestive heart failure, diabetes mellitus with complications, and various gastrointestinal disorders.
+These conditions are known to increase the risk of exacerbations and complications, which are significant predictors of readmission.
+2.
+Procedures: The patient has undergone significant procedures, such as debridement and upper gastrointestinal endoscopy.
+These procedures can lead to complications, necessitating close monitoring and follow-up care, which may contribute to readmission.
+3.
+Medications: The patient is on a complex medication regimen, including beta-blockers and insulins.
+The complexity of this regimen can lead to adherence issues and potential side effects, which might complicate the patient’s health status and increase the likelihood of readmission.
+4.
+Similar Patients: The similar patients, such as Patient ID: 16994_6, experienced complications from chronic diseases and were readmitted.
+This indicates a trend in patients with similar conditions and treatment regimens.
+5.
+Overall Risk Assessment: The combination of chronic conditions, significant procedures, and complex medication regimens suggests a high risk of readmission due to potential exacerbations and complications.
+Given this comprehensive analysis, it is reasonable to conclude that Patient ID: 93900_1 will likely be readmitted to the hospital within 30 days based on the factors outlined above.
+|
+Analysis of distribution (Readmission).
+Figure 3(a) illustrates the distribution of predicted probabilities for the readmission task.
+In the full model, the majority of negatives cluster near zero, while positives are shifted upward and concentrated above a low threshold of 0.16.
+This optimal threshold is chosen using Youden’s J statistic, which maximizes sensitivity and specificity.
+This separation, although not perfectly distinct, reflects the extreme class imbalance in the data: a very low cutoff is required to recover a reasonable fraction of positive cases.
+The overlap between classes explains the modest AUPRC values, since many positives still lie in regions dominated by negatives.
+Nevertheless, compared to ablated variants, the full model achieves tighter grouping of positives in the higher-probability region, resulting in better sensitivity and precision–recall trade-offs.
+The t-SNE plot (Figure 3(b)) further visualizes patients’ features on the held-out test set.
+In this space, readmitted patients do not form sharply separated clusters but instead appear partially embedded within the larger manifold of non-readmitted cases.
+The lack of clear separation reflects the difficulty of the task and the subtlety of signals driving readmission, where positives and negatives overlap substantially.
+Still, there is evidence of localized groupings of readmitted patients, suggesting that the model captures some latent patterns that distinguish higher-risk subgroups.
+This partial clustering is consistent with the modest AUPRC values: while the model cannot fully disentangle the classes, it is able to concentrate a portion of true positives in regions of elevated probability.
+Clinically, this underscores the challenge of predicting readmission but also highlights the value of identifying even partially coherent patient subgroups for targeted follow-up.
+2.2.2 In-Hospital Mortality Prediction
+Mortality prediction is less imbalanced, with approximately positive cases.
+Table 1 reports that KAMELEON-XGBoost and KAMELEON-MLP models achieve high accuracy ( and respectively) and AUROC ( and respectively).
+KAMELEON-XGBoost demonstrates strong performance for mortality prediction, achieving a high precision of , meaning that most patients flagged as high-risk did not survive.
+It also attains a specificity of and an NPV of , indicating that nearly all patients predicted as low-risk were indeed survivors.
+Furthermore, for a positive class that, while less imbalanced, is still a minority, the Area Under the Precision-Recall Curve (AUPRC) is a vital metric.
+Here, the KAMELEON with sets the benchmark, significantly outperforming all other baselines.
+Unstructured models consistently yield the lowest performance for mortality prediction, AUROC values hover just above random chance (around ), and AUPRCs remain very low (max ), indicating a limited ability to discern between mortality and survival solely from clinical notes.
+SHAP results (Appendix Figure 4(b)) demonstrate that lab results and vital signs strongly drive predictions, with prediction embeddings playing a less critical role compared to readmission.
+Overall, Table 1 shows that for both tasks, our multimodal model outperforms all individual structured and unstructured baselines across all metrics.
+Relative importance of different classes of inputs.
+The SHAP values (Figure 4(b)) show that the prediction embedding has a smaller influence compared to labs and vitals, which contrasts with the readmission task.
+The AUROC curve (Figure 4(a)) further illustrates that removing reasoning reduces AUROC from 0.92 in the full model to 0.88, a relative drop of about 4%.
+This indicates that mortality prediction is comparatively easier, as a strong AUROC is retained even without reasoning.
+The lower class imbalance and the strong signal contained in lab values and diagnostic codes allow the model to capture patterns associated with terminal illness more directly.
+Nevertheless, reasoning still contributes by improving discrimination at the margin, capturing subtler risk factors that are less apparent in structured features alone.
+Sample Inference Output from .
+We show inference outputs for two patients on the mortality prediction task, generated by , our finetuned LLM with reasoning in Table 11, Appendix.
+Analysis of distribution (In-Hospital Mortality).
+Figure 5(a) shows predicted probabilities for in-hospital mortality, separated by true class: survivors (blue) and deceased (red).
+The dashed line marks the optimal threshold (0.276) balancing sensitivity and specificity.
+Most survivors cluster near zero probability, reflecting strong model confidence, while deceased cases spread across a wider range, showing prediction uncertainty.
+Overlap between classes causes some misclassifications, highlighting the challenge of predicting this rare event.
+Despite this, the clear separation and tight clustering of survivors demonstrate the model’s strong ability to distinguish between classes, supporting the usefulness of the selected threshold.
+Class Separability Analysis Using t-SNE (In-hospital Mortality Task).
+Figure 5(b) shows a two-dimensional t-SNE embedding of the test data, where the samples are colored by their true class labels (Class 0 in blue and Class 1 in red).
+The visualization reveals that the majority of the data forms a consistent, structured manifold dominated by Class 0 points, with only a small fraction of Class 1 points distributed across the embedding.
+Notably, a compact cluster of Class 1 samples appears on the right-hand side, indicating localized patterns that can be exploited by advanced models.
+This structure suggests that although Class 1 is relatively sparse, it exhibits distinct feature signatures in specific regions, which the proposed KAMELEON‑X model is designed to capture effectively, contributing to improved predictive performance.
+2.2.3 Incorporating additional patient context in
+Our unstructured model, , extends KARE [15] by incorporating physician notes more explicitly into similar patients, enriching the knowledge graph; in contrast, KARE [15] use only structured EHR data related to drugs, procedures, and conditions as context for similar patients, and the knowledge graph is constructed without considering patient conditions.
+Our modification ensures that unstructured clinical narratives contribute to prediction alongside structured features, providing a stronger and more comparable baseline.
+Our strategy for adding physician notes leads to modest improvements in mortality and more substantial gains in readmission (Table 1).
+For mortality prediction, AUROC and AUPRC improve by 2.3% and 1.8%, respectively.
+For readmission, the gains are stronger, with AUROC improving by 3.0% and AUPRC by 2.1%.
+These results suggest that physician notes provide useful complementary signal, especially for readmission, where unstructured narratives capture behavioral, discharge-related risk factors less visible in structured EHR data.
+3 Discussion
+In this work, we introduce KAMELEON, a novel framework that effectively integrates multimodal EHR data, including structured clinical features and unstructured physician notes, enhanced by knowledge-augmented LLM reasoning, for robust clinical risk prediction.
+Our two-stage architecture demonstrated superior performance on both 30-day readmission and in-hospital mortality prediction, with respect to multiple metrics, including the AUROC score.
+KAMELEON outperforms all prior baselines which only used one type of dataset (structured or unstructured), on most metrics for these two tasks, compared to prior work using the MIMIC-III dataset (Table 1).
+Multiple types of standard ML methods have been used for these tasks, yet KAMELEON demonstrates clear improvements across all evaluation metrics.
+None of the currently most powerful LLMs, including a medical LLM trained on clinical data, have comparable performance to KAMELEON.
+The only exception is that the Llama3-Med42-8B model achieves a higher AUPRC for the readmission task; however, KAMELEON significantly outperforms it across all other metrics.
+We find that the reasoning component output by the LLM in , which is used in by constructing an embedding, has high predictive power in both the tasks.
+For the 30-day readmission task, the embedding constructed using the reasoning output by is very significant— removing this component causes the AUROC to drop from to .
+This effect is much smaller in the case of the mortality prediction task, but not negligible, dropping the AUROC from 0.92 to 0.88, when this component is dropped.
+This highlights the synergy achieved by combining these diverse modalities.
+This work underscores the significant potential of knowledge-augmented multimodal EHR modeling to enhance early intervention, optimize resource allocation, and improve patient care in complex clinical settings.
+While LLMs, including medical LLMs trained on specialized data, have a number of limitations in terms of accuracy and hallucinations, their reasoning outputs provide valuable predictive power.
+Future work will focus on further validating KAMELEON’s generalizability across diverse clinical settings and exploring its application to a wider range of predictive healthcare tasks.
+Our framework can be easily extended to other clinical prediction tasks, especially those for which structured models have already been developed.
+KAMELEON can be applied for such tasks without any changes, and we expect it will provide similar gains.
+Scope and Promise for Social Impact.
+KAMELEON offers a strong opportunity to reduce avoidable hospital readmissions, a major driver of morbidity, cost, and financial penalties [41, 42, 43, 44].
+It provides real-time risk predictions for inpatients, enabling more effective discharge planning, case management, and post-acute care.
+Our model uses real-time patient data to assess readmission risk prior to discharge, supporting individualized case management, discharge planning, census forecasting, and post-acute care coordination.
+By identifying high-risk patients, our model enables focused use of limited resources, improving efficiency and outcomes.
+4 Methods
 | Symbol | Description |
@@
 | Task-specific binary cross-entropy loss |
-3 Background
+4.1 Background
 Clinical Datasets. Clinical data used in this study includes both structured and unstructured data collected across patient visits, capturing the patient’s condition over time.
@@
 Problem statements.
-To demonstrate the effectiveness of our method, we study two popular clinical tasks: in-hospital mortality prediction and 30-day readmission prediction, e.g., (Jiang et al.
-2025; Cai et al.
-2016).
+To demonstrate the effectiveness of our method, we study two popular clinical tasks: in-hospital mortality prediction and 30-day readmission prediction, e.g., [15, 1].
 We define these problems formally after introducing some notation.
@@
 Our goal is to build a model that predicts a patient’s target status based on their historical visit information, specifically, .
-30-day Readmission Problem: The objective is to determine whether the patient is readmitted to the hospital within 30 days following discharge from visit . We define the readmission indicator as:
+30-day Readmission Problem. The objective is to determine whether the patient is readmitted to the hospital within 30 days following discharge from visit . We define the readmission indicator as:
 The goal is to develop a predictive model that estimates using all structured and unstructured data from .
@@
 The goal is to develop a predictive model that accurately estimates based on all available structured and unstructured clinical data associated with visit .
-4 KAMELEON Framework
+4.2 KAMELEON Framework
 We propose a hybrid framework, KAMELEON, that integrates multimodal EHR data, including structured clinical components and unstructured physician notes (we will provide more details in Section 2.1), and external biomedical knowledge to predict two key clinical outcomes: in-hospital mortality and 30-day readmission.
-As shown in Figure 1, KAMELEON consists of two components: (1) an unstructured encoder (Section 4.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction, along with its reasoning; and (2) a structured encoder (Section 4.2) that combines multiple time-series corresponding to vitals and tabular datasets (labs, medications, etc), along with the outputs from (i.e., both the prediction and the embedding associated with the reasoning it produces) with static features for downstream prediction.
-The notations used in Section 4, Algorithm 1, and 2 are explained in Table 1.
-4.1 Unstructured Text Encoder,
+As shown in Figure 1, KAMELEON consists of two components: (1) an unstructured encoder (Section 4.2.1) that processes clinical notes and retrieves relevant biomedical knowledge using a PubMed-derived graph and knowledge-augmented reasoning, and outputs a prediction, along with its reasoning; and (2) a structured encoder (Section 4.2.2) that combines multiple time-series corresponding to vitals and tabular datasets (labs, medications, etc), along with the outputs from (i.e., both the prediction and the embedding associated with the reasoning it produces) with static features for downstream prediction.
+The notations used in Section 4.2, Algorithm 1, and 2 are explained in Table 4.
+4.2.1 Unstructured Model,
 For each hospital visit , we collect physician-authored clinical notes and extract entities like conditions, procedures, and medications.
@@
 These triples form a biomedical knowledge graph, serving as an auxiliary source to support LLM reasoning and diagnosis.
-Generating Context: The first step of the framework is generating patient context.
+Generating Context.
+The first step of the framework is generating patient context.
 We use EHR data, including physicians’ notes, patient conditions, prescribed medications, and procedures.
@@
 This approach addresses Challenge 2 (Long Context).
-Knowledge Graph Retrieval: To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)(Bodenreider 2004), PubMed abstracts, and large language model (LLM)-generated entity-relation triples.
+Knowledge Graph Retrieval.
+To introduce domain-level medical knowledge, we build a biomedical knowledge graph (KG) by combining the Unified Medical Language System (UMLS)[29], PubMed abstracts, and large language model (LLM)-generated entity-relation triples.
 UMLS provides standardized biomedical concepts and relationships, and the entity-relation triples are structured facts extracted by the LLM in the form , capturing semantic connections.
-We apply the Leiden algorithm for community detection (Traag, Waltman, and Eck 2019), which partitions the KG into semantically coherent and well-connected subgraphs.
+We apply the Leiden algorithm for community detection [45], which partitions the KG into semantically coherent and well-connected subgraphs.
 After clustering, we employ a separate LLM to generate a textual summary for each cluster.
@@
 Each summary serves as a high-level abstraction of the biomedical concepts and interactions within a subgraph.
-We embed these summaries using SentenceTransformer (MiniLM-L6-v2) (Reimers and Gurevych 2019) and retrieve the most relevant ones for each patient by computing semantic similarity with the embedded patient context.
+We embed these summaries using SentenceTransformer (MiniLM-L6-v2) [46] and retrieve the most relevant ones for each patient by computing semantic similarity with the embedded patient context.
 This process directly addresses Challenge 3 (Specialized Medical Domains) by enriching patient context with structured, domain-specific knowledge, improving the model’s understanding of specialized medical terminology.
-Finding Similar Patients: We provide additional context, by retrieving semantically similar patient visits using precomputed visit-level embeddings and a similarity index using the FAISS (Douze et al.
-2024) library with inner-product search on L2-normalized embeddings, which effectively approximates cosine similarity.
+Finding Similar Patients.
+We provide additional context, by retrieving semantically similar patient visits using precomputed visit-level embeddings and a similarity index using the FAISS [47] library with inner-product search on L2-normalized embeddings, which effectively approximates cosine similarity.
 For each target patient visit, we retrieve the top 50 most similar patients while excluding self-matches and other visits from the same individual.
@@
 The final output includes the top-k positive and negative similar patients (with k=1,2).
-Unlike KARE (Jiang et al.
-2025), we also provide the physician notes of the retrieved similar patients, enabling the language model to leverage more clinical context when assessing patient risk.
-Reasoning Module: In this module, we prepare inputs to fine-tune the LLM for clinical prediction.
+Unlike KARE [15], we also provide the physician notes of the retrieved similar patients, enabling the language model to leverage more clinical context when assessing patient risk.
+Reasoning Module.
+In this module, we prepare inputs to fine-tune the LLM for clinical prediction.
 For each patient visit, we create a prompt with the patient’s context with the top-k most similar cases retrieved earlier.
@@
 Combining patient data, historical cases, and domain knowledge, we fine-tune the LLM to produce task-specific predictions with interpretable reasoning, supporting each outcome.
-Finetuning LLM: We fine-tune a LLaMA-3 8B model using the Unsloth framework (Touvron et al.
-2024; Daniel Han and team 2023), which enables memory-efficient training via 4-bit quantization and Low-Rank Adaptation (LoRA) (Hu et al.
-2022).
+Finetuning LLM.
+We fine-tune a LLaMA-3 8B model using the Unsloth framework [48, 49], which enables memory-efficient training via 4-bit quantization and Low-Rank Adaptation (LoRA) [50].
 Prediction tasks are framed as instruction-following using Alpaca-style prompts with task description, patient context, and optional justification.
@@
 Additionally, an example prompt and its overall structure are provided in the Appendix Section A.3.
-4.2 Structured Data Encoder,
+4.2.2 Structured Data Encoder,
 We extract structured features from the patient’s visit history during each stay, including:
-Time-varying variables: We extract hourly vitals and diagnoses in MIMIC-III, including heart rate, systolic and diastolic blood pressure, mean blood pressure, oxygen saturation, Glasgow Coma Scale (GCS) scores, glucose level, respiratory rate, temperature, weight, and pH.
-Static metadata: Demographic and admission features include gender, ethnicity, admission type, location, insurance, language, religion, and hospital outcome.
-Diagnoses, procedures, medications: ICD-9 codes, drug names are encoded via one-hot or counts.
+Time-varying variables.
+We extract hourly vitals and diagnoses in MIMIC-III, including heart rate, systolic and diastolic blood pressure, mean blood pressure, oxygen saturation, Glasgow Coma Scale (GCS) scores, glucose level, respiratory rate, temperature, weight, and pH.
+Static metadata. Demographic and admission features include gender, ethnicity, admission type, location, insurance, language, religion, and hospital outcome.
+Diagnoses, procedures, medications.
+ICD-9 codes, drug names are encoded via one-hot or counts.
 We compute binary indicators for key comorbidities (e.g., sepsis, infection, organ failure, dementia, cancer, diabetes).
-Structured Data Preprocessing: The structured data is first transformed using a discretization step to enforce uniform temporal resolution and impute missing values.
+Structured Data Preprocessing.
+The structured data is first transformed using a discretization step to enforce uniform temporal resolution and impute missing values.
 This is followed by normalization using precomputed mean and standard deviation statistics over the continuous variables.
-Incorporating LLM Output: To augment the structured input, we include the LLM’s prediction and its tokenized reasoning.
-A Word2Vec (Mikolov et al.
-2013) model is trained on this reasoning to generate embeddings.
+Incorporating LLM Output.
+To augment the structured input, we include the LLM’s prediction and its tokenized reasoning.
+A Word2Vec [51] model is trained on this reasoning to generate embeddings.
 Each paragraph is represented by averaging its word vectors, producing a fixed-size vector.
@@
 To reduce dimensionality and suppress noise from high-dimensional embeddings, we apply Principal Component Analysis (PCA) to the combined feature vector.
-4.3 Training
+4.2.3 Training
 We follow a two-stage training procedure.
@@
 After fine-tuning, we perform Final Integration and use the outputs of as input features to train for final prediction.
-In our experiments, we benchmark several ML models for as KAMELEON-X, where X represents logistic regression, random forests, LSTM, LightGBM, MLP, or XGBoost, selected for its effectiveness in capturing clinical patterns.
+In our experiments, we benchmark several ML models for as KAMELEON-X, where X represents logistic regression, balanced random forests (BRF), LSTM, LightGBM, MLP, or XGBoost, selected for its effectiveness in capturing clinical patterns.
 For MLP, we use weighted binary cross-entropy loss
 where and are the positive and negative class weights, respectively, used to address class imbalance by giving more emphasis to the minority class.
-We further use Synthetic Minority Over-sampling Technique (SMOTE) (Chawla et al.
-2002) to mitigate class imbalance, addressing Challenge 4 (Highly imbalanced data).
-The pipeline for is in Algorithm 2.
-| Model | Acc | Prec0 | Prec1 | Rec0 | Rec1 | Macro F1 | AUROC | AUPRC | ||
-| Task: 30-Day Readmission Prediction | ||||||||||
-| LogReg (C=1.0) (Harutyunyan et al. 2019) | ✓ | – | 0.831 | 0.869 | 0.036 | 0.951 | 0.013 | 0.464 | 0.463 | 0.090 |
-| MLP (Bopche et al. 2024) | ✓ | – | 0.828 | 0.876 | 0.182 | 0.934 | 0.100 | 0.516 | 0.559 | 0.165 |
-| BalancedRF | ✓ | – | 0.760 | 0.970 | 0.070 | 0.780 | 0.430 | 0.490 | 0.673 | 0.066 |
-| LSTM (Harutyunyan et al. 2019) | ✓ | – | 0.820 | 0.876 | 0.163 | 0.925 | 0.100 | 0.512 | 0.569 | 0.152 |
-| Claude-3.7-Sonnet | – | ✓ | 0.240 | 0.790 | 0.199 | 0.068 | 0.927 | 0.227 | 0.498 | 0.199 |
-| Medgemma-4b-it | – | ✓ | 0.350 | 0.770 | 0.190 | 0.270 | 0.690 | 0.350 | 0.480 | 0.190 |
-| Llama3-Med42-8B | – | ✓ | 0.390 | 0.800 | 0.210 | 0.360 | 0.670 | 0.410 | 0.510 | 0.210 |
-| KARE | – | ✓ | 0.660 | 0.870 | 0.130 | 0.720 | 0.280 | 0.480 | 0.506 | 0.195 |
-| KAMELEON-LogReg, (C=1.0) | ✓ | ✓ | 0.833 | 0.869 | 0.037 | 0.953 | 0.013 | 0.519 | 0.130 | 0.148 |
-| KAMELEON-LogReg, (C=0.01) | ✓ | ✓ | 0.871 | 0.874 | 0.333 | 0.996 | 0.013 | 0.478 | 0.551 | 0.152 |
-| KAMELEON-LSTM | ✓ | ✓ | 0.840 | 0.880 | 0.190 | 0.950 | 0.090 | 0.510 | 0.505 | 0.135 |
-| KAMELEON-BalancedRF | ✓ | ✓ | 0.800 | 0.990 | 0.130 | 0.800 | 0.790 | 0.550 | 0.845 | 0.150 |
-| KAMELEON-MLP | ✓ | ✓ | 0.910 | 0.970 | 0.160 | 0.940 | 0.280 | 0.580 | 0.820 | 0.138 |
-| Task: In-Hospital Mortality Prediction | ||||||||||
-| Logistic Regression (Harutyunyan et al. 2019) | ✓ | – | 0.850 | 0.912 | 0.340 | 0.916 | 0.331 | 0.625 | 0.624 | 0.190 |
-| LSTM (Harutyunyan et al. 2019) | ✓ | – | 0.690 | 0.800 | 0.260 | 0.800 | 0.250 | 0.530 | 0.560 | 0.240 |
-| BalancedRF | ✓ | – | 0.810 | 0.950 | 0.340 | 0.820 | 0.700 | 0.670 | 0.860 | 0.475 |
-| LightGBM (Bopche et al. 2024) | ✓ | – | 0.890 | 0.930 | 0.510 | 0.940 | 0.480 | 0.720 | 0.866 | 0.534 |
-| MLP (Bopche et al. 2024) | ✓ | – | 0.870 | 0.920 | 0.430 | 0.920 | 0.430 | 0.680 | 0.829 | 0.426 |
-| XGBoost | ✓ | – | 0.890 | 0.920 | 0.520 | 0.950 | 0.380 | 0.695 | 0.835 | 0.487 |
-| Claude-3.7-Sonnet | – | ✓ | 0.800 | 0.890 | 0.120 | 0.880 | 0.130 | 0.510 | 0.510 | 0.110 |
-| Medgemma-4b-it | – | ✓ | 0.120 | 0.950 | 0.100 | 0.020 | 0.990 | 0.120 | 0.510 | 0.110 |
-| Llama3-Med42-8B | – | ✓ | 0.160 | 0.950 | 0.120 | 0.100 | 0.970 | 0.190 | 0.530 | 0.120 |
-| KARE | – | ✓ | 0.614 | 0.890 | 0.134 | 0.641 | 0.413 | 0.474 | 0.527 | 0.125 |
-| KAMELEON-BalancedRF | ✓ | ✓ | 0.880 | 0.930 | 0.490 | 0.934 | 0.492 | 0.710 | 0.876 | 0.543 |
-| KAMELEON-LSTM | ✓ | ✓ | 0.730 | 0.820 | 0.430 | 0.840 | 0.390 | 0.620 | 0.740 | 0.350 |
-| KAMELEON-lightGBM | ✓ | ✓ | 0.880 | 0.940 | 0.470 | 0.910 | 0.590 | 0.730 | 0.890 | 0.550 |
-| KAMELEON-MLP | ✓ | ✓ | 0.900 | 0.940 | 0.550 | 0.940 | 0.550 | 0.750 | 0.890 | 0.600 |
-| KAMELEON-XGBoost | ✓ | ✓ | 0.920 | 0.920 | 0.790 | 0.980 | 0.369 | 0.660 | 0.920 | 0.650 |
-5 Results
-We evaluate our model on the two clinical prediction tasks: in-hospital mortality and 30-day readmission.
-Our experiments compare performance against strong baselines, including general and medical LLMs, and assess the value of key model components through ablation studies.
-Some of our results are discussed in Section B Appendix.
-5.1 Experimental Setup
-Datasets.
-We use the MIMIC-III dataset (Johnson et al.
-2016), which includes structured and unstructured data for over 40,000 ICU patients.
-It includes structured data (demographics, vitals, labs, admissions, ICD-9-CM codes) and unstructured clinical text (physician notes, discharge summaries, radiology reports).
-For this study, we focus on physician-authored notes containing clinical reasoning, assessments, and treatment plans.
-Here, we exclude discharge summaries and notes written after outcomes to prevent label leakage.
-Only of notes mention hospice, indicating rare explicit terminal indicators.
-Biomedical Knowledge.
-We use abstracts from the annual PubMed Baseline dataset, comprising over 36 million biomedical citation records, to build a medical knowledge graph that enriches LLM input, and reduces hallucinations.
-We also incorporate UMLS (Bodenreider 2004) to construct concept-centric subgraphs from EHR data.
-Dataset Statistics.
-We include a summary table in the Appendix, showing dataset statistics, indicatin moderate class imbalance for in-hospital mortality (positive) and severe imbalance for 30-day readmission ( positive).
-To prevent data leakage, train-test splits are patient disjoint, meaning that multiple visits from the same patient do not appear in both sets.
-Baselines.
-We compare against Claude 3.5 Sonnet (Anthropic 2024), MedGemma (Sellergren et al.
-2025), LLaMA3-Med (Christophe et al.
-2024), and KARE (Jiang et al.
-2025), as well as structured-data models including logistic regression, tree-based models, and MLPs used in prior work on MIMIC III (Harutyunyan et al.
-2019; Lundberg et al.
-2020; Purushotham et al.
-2018).
-All LLMs are evaluated in a zero-shot setting with the same patient-context prompt.
-KARE uses similar patient retrieval but lacks clinical notes.
-Our model incorporates retrieved notes for better context.
-Implementation details are in the Appendix.
-Metrics. We report AUROC, AUPRC, accuracy, precision, recall, and macro F1-score. Metrics are computed on a held-out test set with similar standardized preprocessing.
-5.2 Performance of KAMELEON
-30-Day Readmission Prediction:
-Readmission within 30 days is a highly imbalanced task, with only about positive cases in the dataset.
-This severe imbalance is reflected in the results, where most models achieve high accuracy and precision on the negative class but struggle with recall for the positive (readmitted) class.
-As shown in Table 2, our framework with Balanced Random Forest (KAMELEON-BalancedRF) classifier achieves the highest AUROC and notably improves recall on positive cases to , a crucial metric since identifying patients at risk of readmission is clinically imperative.
-The KAMELEON-MLP model, while achieving the highest overall accuracy and macro F1 , still attains a recall of on positive cases, illustrating the persistent challenge in detecting rare events.
-Unstructured LLM-based baselines such as Claude-3.7-Sonnet, MedGemma, LLaMA3-Med, and KARE show substantially lower recall for positives (below 0.3), suggesting that these models struggle to identify the minority class without further fine-tuning or domain-specific adaptation.
-To better understand this model’s behavior, we perform SHAP analysis to identify feature importance, and Figure 2 indicates that the model relies primarily on prediction embeddings and lab/vital features for predicting 30-day readmission, highlighting the importance of multimodal inputs.
-While prior studies like Morgan et al.
-(2019) (Morgan et al.
-2019) reported AUROCs up to 0.81 for readmission, and general models typically ranged from 0.61 to 0.73 (Dhalluin et al.
-2020; Matheny et al.
-2021), our multi-modal approach effectively captures complex clinical nuances.
-In-Hospital Mortality Prediction: Mortality prediction is less imbalanced, with approximately positive cases.
-Table 2 reports that KAMELEON-XGBoost and KAMELEON-MLP models achieve high accuracy ( and respectively) and AUROC ( and respectively).
-Furthermore, for a positive class that, while less imbalanced, is still a minority, the Area Under the Precision-Recall Curve (AUPRC) is a vital metric.
-Here, the KAMELEON with sets the benchmark, significantly outperforming all other baselines.
-Unstructured models consistently yield the lowest performance for mortality prediction, AUROC values hover just above random chance (around ), and AUPRCs remain very low (max ), indicating a limited ability to discern between mortality and survival solely from clinical notes.
-SHAP results (Appendix Figure 8) demonstrate that lab results and vital signs strongly drive predictions, with prediction embeddings playing a less critical role compared to readmission.
-Overall, Table 2 shows that for both tasks, our multimodal model outperforms all individual structured and unstructured baselines across all metrics.
-5.3 Ablation Study
-To highlight the impact of reasoning features, we perform an ablation study (Table 8, Appendix).
-Removing reasoning features alone causes AUROC to drop from to for 30-day readmission.
-When we further remove demographics and prediction embeddings alongside reasoning, performance falls below , indicating that each component plays a complementary role.
-Figure 2 visualizes the contribution of reasoning, ’s prediction, and demographics to model performance.
-We also provide a similar plot for in-hospital mortality prediction in the Appendix.
-For this task, structured data played a more critical role, with less relative impact from reasoning and unstructured features compared to readmission prediction.
-6 Conclusion
-In this work, we introduced KAMELEON, a novel framework that effectively integrates multimodal EHR data, including structured clinical features and unstructured physician notes, enhanced by knowledge-augmented LLM reasoning, for robust clinical risk prediction.
-Our two-stage architecture demonstrated superior performance on both 30-day readmission and in-hospital mortality prediction, achieving a high AUROC score despite severely imbalanced datasets.
-KAMELEON consistently outperformed existing models, highlighting the synergy achieved by combining these diverse modalities.
-This work underscores the significant potential of knowledge-augmented multimodal EHR modeling to enhance early intervention, optimize resource allocation, and improve patient care in complex clinical settings.
-Future work will focus on further validating KAMELEON’s generalizability across diverse clinical settings and exploring its application to a wider range of predictive healthcare tasks.
-Scope and Promise for Social Impact.
-KAMELEON offers a strong opportunity to reduce avoidable hospital readmissions, a major driver of morbidity, cost, and financial penalties (Fry, Fluck, and Han 2021; Joynt and Jha 2013; Panagiotou et al.
-2021; Zuckerman et al.
-2017).
-The greatest impact is expected in high-risk groups, such as Outpatient Parenteral Antimicrobial Therapy (OPAT) patients, who face up to readmission rates (Wolie et al.
-2024).
-OPAT programs are often underfunded and rely on manual, error-prone reviews.
-By identifying high-risk patients, our model enables focused use of limited resources, improving efficiency and outcomes.
-We are exploring deployment in a large regional hospital for such high-impact populations, with later expansion to general readmissions.
+We further use Synthetic Minority Over-sampling Technique (SMOTE) [52] to mitigate class imbalance, addressing Challenge 4 (Highly imbalanced data).
+For KAMELEON-BRF model, which is designed to handle class imbalance.
+Here, each decision tree is trained on a bootstrapped sample drawn by under-sampling the majority class and combining it with all minority class examples, ensuring balanced class proportions at the tree level.
+Each split is chosen to minimize the Gini impurity:
+where denotes teh proportion of class in a given node.
+The balanced sampling overcomes the issue with the dominance of the majority class and improves sensitivity to rare outcomes.
+Final predictions are obtained by aggregating probabilities across trees.
+The complete training and inference pipeline for is outlined in Algorithm 2.
 Appendix A Additional Details on Dataset
-We use the Medical Information Mart for Intensive Care III (MIMIC-III) dataset (Johnson et al.
-2016), a large-scale, publicly available critical care database comprising deidentified health-related data associated with over patients who were admitted to the intensive care units (ICUs) of Beth Israel Deaconess Medical Center between and .
+We use the Medical Information Mart for Intensive Care III (MIMIC-III) dataset [35], a large-scale, publicly available critical care database comprising deidentified health-related data associated with over patients who were admitted to the intensive care units (ICUs) of Beth Israel Deaconess Medical Center between and .
 The dataset was curated to support clinical and epidemiological research and has since become a cornerstone resource in machine learning for healthcare.
@@
 In the main paper, we described multiple stages of unstructured text encoder, . In this section, we will provide, examples on each step.
-Generating Context
+A.3.1 Generating Context
 In this section, for each patient ’s visit , we utilize information from the MIMIC-III dataset, including ICD codes and physician notes.
@@
 Figure 6 demonstrates a sample example of a patient context.
-Knowledge Graph Retrieval
+A.3.2 Knowledge Graph Retrieval
 The first step in our approach involves downloading the latest PubMed dataset and extracting all available abstracts.
@@
 KG structure: This PubMed-derived knowledge graph supplements patient-specific data and improves reasoning by providing broader clinical context.
-The following triples represent extracted relations involving beta blocking agents from biomedical literature:
+The following triples are an example how the relations are extracted, involving beta blocking agents from biomedical literature.
 (unstable angina pectoris, treated with, beta blocking agents)
@@
 (beta blocking agents, reduce, arterial NEFA levels)
-Finding Similar Patients
+A.3.3 Finding Similar Patients
 Here, the model identifies similar patients by leveraging diagnoses, procedures, and medications, along with unstructured data from physician notes.
@@
 | Label | 1 | 0 |
-Reasoning Module
+A.3.4 Reasoning Module
 A critical component of our framework is the reasoning module, which enables the model to go beyond surface-level pattern matching and perform clinically informed inference.
@@
 The prompt concludes with an explicit request for medically grounded reasoning followed by the final risk prediction. This structured input design facilitates interpretable and accurate model outputs.
-Sample Inference Output
-Table 7 shows a sample output generated by the fine-tuned LLM for the 30-day hospital readmission prediction task.
-Instead of returning a raw binary value (0 or 1), the model is prompted to generate both a prediction and its reasoning based on the patient’s diagnoses, procedures, medications, and comparisons with similar cases.
-This structured explanation allows the model to ground its prediction in clinical context, improving reliability and interpretability over naive classification.
-By incorporating rationale into the output, the LLM demonstrates better alignment with real-world clinical decision-making.
-Patient ID: 25070_0
-To predict whether Patient ID 25070_0 will be readmitted within 30 days, we analyze the provided EHR context, focusing on conditions, procedures, and medications.
-1.
-Conditions: The patient has multiple serious conditions, including non-Hodgkin’s lymphoma and complications from a device, which are significant risk factors for readmission.
-However, they also have other conditions that may not be as acute (e.g., asthma, thyroid disorders).
-2.
-Procedures: The patient has undergone several diagnostic procedures, including mastoidectomy, which could lead to complications but are not major surgeries.
-The patient did not have any major surgical interventions that typically increase readmission risk.
-3.
-Medications: The patient is on a complex regimen including corticosteroids and other systemic medications.
-The presence of corticosteroids indicates a potential for complications, particularly in the management of their non-Hodgkin’s lymphoma.
-4.
-Similar Patients: The similar patients show that those with chronic obstructive pulmonary disease and heart failure have higher readmission rates.
-However, Patient ID 25070_0 does not share the same chronic conditions with these patients.
-5.
-Overall Assessment: Although the patient has several chronic and acute conditions, they do not have the same severity or complexity as the patients who have been readmitted.
-The absence of major recent surgeries and the complexity of their medication regimen suggests a lower risk of readmission.
-# Prediction # 0 |
-Patient ID: 93900_1
-To predict whether Patient ID: 93900_1 will be readmitted to the hospital within 30 days, we analyze the patient’s EHR context, focusing on conditions, procedures, and medications.
-1.
-Conditions: The patient has multiple chronic conditions, including congestive heart failure, diabetes mellitus with complications, and various gastrointestinal disorders.
-These conditions are known to increase the risk of exacerbations and complications, which are significant predictors of readmission.
-2.
-Procedures: The patient has undergone significant procedures, such as debridement and upper gastrointestinal endoscopy.
-These procedures can lead to complications, necessitating close monitoring and follow-up care, which may contribute to readmission.
-3.
-Medications: The patient is on a complex medication regimen, including beta-blockers and insulins.
-The complexity of this regimen can lead to adherence issues and potential side effects, which might complicate the patient’s health status and increase the likelihood of readmission.
-4.
-Similar Patients: The similar patients, such as Patient ID: 16994_6, experienced complications from chronic diseases and were readmitted.
-This indicates a trend in patients with similar conditions and treatment regimens.
-5.
-Overall Risk Assessment: The combination of chronic conditions, significant procedures, and complex medication regimens suggests a high risk of readmission due to potential exacerbations and complications.
-Given this comprehensive analysis, it is reasonable to conclude that Patient ID: 93900_1 will likely be readmitted to the hospital within 30 days based on the factors outlined above.
-|
 A.5 Structured Data Representation for
@@
 Appendix B Additional results
-Ablation on Readmission Table 8 presents an ablation study highlighting the performance impact of removing specific input features and components from models and for the readmission prediction task.
-The variations include removal of reasoning components, demonstrations, and predicted features.
-KAMELEON achieves an accuracy of 0.80 and a Macro F1 score of 0.55, indicating a balanced performance across both classes.
-When the reasoning component is removed, we observe a spike in accuracy (0.92), but a sharp drop in recall for the positive class (Rec1 = 0.13), showing that the model becomes overly biased toward the negative class.
-This is further reflected in the reduced AUPRC and AUROC values.
-Removing reasoning, demonstrations, and predicted features together leads to the worst performance (Macro F1 = 0.39), confirming that these components are critical for model generalization, especially in correctly identifying positive cases.
-The stepwise removal of features helps isolate the contribution of each component, underlining the importance of structured reasoning and prior knowledge in our approach.
-Ablation on Mortality However, the SHAP values (Figure 8) show that the prediction embedding has a smaller impact compared to lab vitals, which contrasts with the readmission task.
-Additionally, we observe a higher AUROC for in-hospital mortality even without reasoning, exceeding .
-This high AUROC score suggests that predicting mortality is comparatively easier, likely due to the dataset being less imbalanced.
-Given access to lab vitals and diagnostic information, the model can more easily learn patterns associated with terminal illness.
-| Model | Acc | Prec0 | Prec1 | Rec0 | Rec1 | Macro F1 | AUROC | AUPRC |
-| our setup | 0.80 | 0.99 | 0.13 | 0.80 | 0.77 | 0.55 | 0.844 | 0.147 |
-| without reasoning | 0.92 | 0.97 | 0.11 | 0.96 | 0.13 | 0.54 | 0.687 | 0.082 |
-| without demo, reasoning, pred | 0.54 | 0.97 | 0.05 | 0.54 | 0.58 | 0.39 | 0.559 | 0.0433 |
-| without reasoning, pred | 0.57 | 0.97 | 0.05 | 0.57 | 0.62 | 0.41 | 0.594 | 0.048 |
-Analysis of distribution (In-Hospital Mortality)
-Figure 9 shows predicted probabilities for in-hospital mortality within 30 days, separated by true class: survivors (blue) and deceased (red).
-The dashed line marks the optimal threshold (0.26) balancing sensitivity and specificity.
-Most survivors cluster near zero probability, reflecting strong model confidence, while deceased cases spread across a wider range, showing prediction uncertainty.
-Overlap between classes causes some misclassifications, highlighting the challenge of predicting this rare event.
-Despite this, the clear separation and tight clustering of survivors demonstrate the model’s strong ability to distinguish between classes, supporting the usefulness of the selected threshold.
-Class Separability Analysis Using t-SNE (In-hospital Mortality Task)
-Figure 10 shows a two-dimensional t-SNE embedding of the test data, where the samples are colored by their true class labels (Class 0 in blue and Class 1 in red).
-The visualization reveals that the majority of the data forms a consistent, structured manifold dominated by Class 0 points, with only a small fraction of Class 1 points distributed across the embedding.
-Notably, a compact cluster of Class 1 samples appears on the right-hand side, indicating localized patterns that can be exploited by advanced models.
-This structure suggests that although Class 1 is relatively sparse, it exhibits distinct feature signatures in specific regions, which the proposed KAMELEON‑X model is designed to capture effectively, contributing to improved predictive performance.
 Using Anomaly Detection Method for Readmission in 30 days: Given the highly imbalanced nature of the 30-day readmission data, where only approximately of hospital visits result in readmission, we explored whether this task could be framed as an anomaly detection problem.
@@
 | Test Set | 123 | 3113 | 130 | 3 |
+Patient ID: 9782_1
+### Step 1: Analyzing Conditions 1. **Septicemia (except in labor)** - This is a severe condition that indicates systemic infection, which can lead to sepsis and high mortality risk. |
+| 2. **Shock** - Indicates a critical state, often associated with poor outcomes. |
+| 3. **Pneumonia** - This can be serious, especially in the context of septicemia and shock, as it can complicate respiratory function. |
+| 4. **Schizophrenia and other psychotic disorders** - While this is a mental health condition, it does not directly impact mortality. |
+| 5. **Deficiency and other anemia** - This can complicate other conditions, especially in the context of infection or respiratory issues. |
+| 6. **Thyroid disorders** - Generally less critical unless there are acute complications. |
+| ### Step 2: Analyzing Procedures 1. **Other vascular catheterization; not heart** - This is an invasive procedure that can carry risks, especially in the context of the patient’s other conditions. |
+| 2. **No major surgeries or life-sustaining procedures** - The absence of life-sustaining interventions suggests that the patient may not be in immediate critical condition. |
+| ### Step 3: Analyzing Medications 1. **Antibiotics (quinolone antibacterials, other antibacterials)** - Indicate the treatment of infection, which is critical given the septicemia and pneumonia. |
+| 2. **Corticosteroids for systemic use** - Can be used in severe conditions but also carry risks, especially in the presence of infections. |
+| 3. **Opioid analgesics** - Often used for pain management, which may indicate severe underlying conditions. |
+| 4. **Cardiac stimulants** - Suggests potential heart issues. |
+| 5. **High-risk medications** - The use of multiple medications indicates a complex health status, which increases the risk of adverse effects and complications. |
+| ### Step 4: Comparing with Similar Patients 1.
+**Patient ID 15209_0** had conditions such as septicemia, respiratory failure, and shock, leading to a label of 0 (survival).
+This patient had similar severe conditions but also underwent extensive invasive procedures and had more acute conditions.
+|
+| 2.
+**Patient ID 28061_0** had coronary atherosclerosis and cardiac issues leading to a label of 1 (mortality).
+This patient had significant cardiovascular issues and underwent multiple invasive procedures.
+|
+| ### Step 5: Integrating Information |
+| - The combination of septicemia, shock, and pneumonia in Patient ID 9782_1 is extremely concerning, especially considering the lack of major surgical interventions.
+The presence of multiple comorbidities (anemia, schizophrenia, and thyroid disorders) complicates the overall picture.
+|
+| - The complexity and severity of the patient’s conditions (particularly septicemia and shock) indicate a high risk of mortality, but the absence of life-sustaining procedures suggests some resilience.
+|
+| ### Conclusion |
+| Given the severity and complexity of Patient ID 9782_1’s conditions, the presence of multiple high-risk medications, and the lack of immediate life-threatening interventions, the prediction leans towards survival (label 0).
+However, the critical nature of the patient’s conditions suggests that the risk of mortality is significant, and a careful monitoring plan is essential for managing the patient’s health.
+|
+| ### Final Prediction |
+| - **Prediction**: 0 (survival) |
+Patient ID: 85258_1
+### Step 1: Analyze Patient Conditions |
+| 1. **Cancer of Esophagus**: This is a critical condition that is often associated with high mortality risk, particularly when combined with other severe conditions. |
+| 2. **Secondary Malignancies**: The presence of secondary malignancies indicates that the patient has advanced cancer, further increasing the risk of mortality. |
+| 3. **Pneumonia**: This is a severe infection that can lead to respiratory failure, especially in a patient with compromised respiratory function due to cancer. |
+| 4. **Respiratory Failure**: The patient has been diagnosed with respiratory failure, indicating a critical state that significantly increases the risk of mortality. |
+| 5. **Septicemia**: This is a serious systemic infection that can lead to multi-organ failure, which is a key indicator of mortality. |
+| 6. **Deficiency Anemia**: This can complicate the patient’s overall health, particularly in the context of cancer and infection. |
+| 7.
+**Other Comorbidities**: The patient has several other comorbidities, including disorders of teeth and jaw, white blood cell diseases, and fluid and electrolyte disorders, which compound the risk of mortality.
+|
+| ### Step 2: Evaluate Procedures |
+| 1. **Tracheostomy**: This is a significant surgical intervention that indicates severe respiratory compromise and is often performed in critically ill patients. |
+| 2. **Mechanical Ventilation**: The patient is on mechanical ventilation, which is a life-sustaining measure typically reserved for critically ill patients. |
+| 3. **Blood Transfusions**: The need for multiple blood transfusions suggests significant anemia or other severe conditions that require urgent intervention. |
+| ### Step 3: Assess Medications |
+| 1. **Broad Spectrum Antibiotics**: The patient is receiving broad-spectrum antibiotics, which are often used in the presence of sepsis or severe infections. |
+| 2. **Corticosteroids**: The use of corticosteroids suggests a significant inflammatory response or severe underlying condition. |
+| 3. **Antipsychotics and Antidepressants**: These medications suggest a complex mental health component, which can complicate treatment and recovery. |
+| ### Step 4: Compare with Similar Patients |
+| 1.
+**Patient ID: 28061_0**: This patient had a history of coronary artery disease and complications from medical care, leading to a mortality label of 1.
+The presence of critical cardiac issues and multiple comorbidities mirrors the complexity of Patient ID: 85258_1.
+|
+| 2.
+**Patient ID: 14297_0**: This patient had several chronic conditions but ultimately had a label of 0, indicating that not all patients with complex conditions lead to mortality.
+The complexity of Patient ID: 85258_1, however, suggests a much higher risk.
+|
+| ### Step 5: Synthesize Findings |
+| - The combination of advanced cancer, severe infections (septicemia), respiratory failure, and invasive procedures (tracheostomy and mechanical ventilation) strongly indicates that the patient is in a critical state.
+|
+| - The presence of multiple comorbidities and the complexity of the medications suggest that the patient is in a high-risk category for mortality. |
+| ### Conclusion |
+| Given the severity of the conditions, the invasive nature of procedures, and the complex medication regimen, the patient is at an extremely high risk of mortality.
+Therefore, the prediction for Patient ID: 85258_1 is: |
+| **Prediction**: 1 (mortality) |
+Appendix C Related Works
+Clinical Prediction using Statistical and Machine Learning Methods: Structured data collected from electronic health records (EHRs) contain valuable information about patients and can be instrumental in capturing patient states and predicting medical outcomes (e.g., mortality) [15, 16, 17].
+To leverage these data, traditional machine learning models—such as linear regression [8], decision trees [9], and Bayesian approaches [18]—use structured features as input to model the complex relationships between patient features and clinical outcomes.
+More recently, deep learning–based frameworks have also been explored to capture these complex dynamics and have shown great potential [10, 11, 12, 13, 14].
+However, deep learning-based approaches remain largely restricted to structured data and rely heavily on statistical correlations between input features and outcome labels.
+Deep learning-based approaches also often lack interpretability grounded in explicit clinical knowledge and tend to be highly specific to the training dataset, limiting their generalizability to other clinical settings [15, 26].
+NLP and Language Models: Unstructured clinical texts (e.g., physician notes, radiology reports, and nursing progress notes) contain rich information about patients [35, 53, 26].
+Textual documentation includes clinicians’ reasoning and management plans that are not contained elsewhere in the EHR and which may provide direct and high-value insights.
+To harness this information, various natural language processing (NLP) methods have been proposed for predictive modeling [54, 55, 56].
+Recently, large language models (LLMs) have demonstrated significant progress across a range of tasks, and researchers have begun exploring their potential for processing unstructured clinical text and supporting clinical reasoning [53, 28, 27].
+However, one of the major challenges in applying LLMs to healthcare is the risk of errors and hallucinations [57, 22].
+To address this, some works incorporate knowledge graphs to explicitly embed clinical knowledge and mitigate LLM inaccuracies [58, 26].
+Despite these advances, integrating both structured EHR data and unstructured clinical text in a unified framework remains a challenging problem [59].
+Existing approaches—such as prompt engineering [60], RAG [14], and data fusion [61, 62]—tend to be ad hoc and lack a systematic method for joint modeling.